### v2

# A LoD of Gaussians: Unified Training and Rendering for Ultra-Large Scale Reconstruction with External Memory

Abstract.
Gaussian Splatting has emerged as a high-performance technique for novel view synthesis, enabling real-time rendering and high-quality reconstruction of small scenes. However, scaling to larger environments has so far relied on partitioning the scene into chunks—a strategy that introduces artifacts at chunk boundaries, complicates training across varying scales, and is poorly suited to unstructured scenarios such as city-scale flyovers combined with street-level views. Moreover, rendering remains fundamentally limited by GPU memory, as all visible chunks must reside in VRAM simultaneously. We introduce A LoD of Gaussians, a framework for training and rendering ultra-large-scale Gaussian scenes on a single consumer-grade GPU—without partitioning. Our method stores the full scene out-of-core (e.g., in CPU memory) and trains a Level-of-Detail (LoD) representation directly, dynamically streaming only the relevant Gaussians. A hybrid data structure combining Gaussian hierarchies with Sequential Point Trees enables efficient, view-dependent LoD selection, while a lightweight caching and view scheduling system exploits temporal coherence to support real-time streaming and rendering. Together, these innovations enable seamless multi-scale reconstruction and interactive visualization of complex scenes—from broad aerial views to fine-grained ground-level details.
Level of Detail, Gaussian Splatting, Large-Scale Reconstruction
isbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Rasterizationccs: Computing methodologies Visibilityccs: Computing methodologies Machine learning
Figure 1. We introduce a fully hierarchical 3D Gaussian representation trained directly across unstructured, multi-scale image sets—including street-level and aerial views—without scene partitioning. Our method maintains a consistent global scene model, eliminating boundary artifacts typical of chunked approaches. A hybrid Level-of-Detail system combines Gaussian hierarchies with Sequential Point Trees, enabling dynamic, view-dependent streaming and LoD selection. The entire model resides in external memory, with only a small, adaptive subset loaded on demand, allowing real-time rendering and training of 60M+ Gaussians on consumer GPUs ($\leq 24GB$ VRAM). This is the first approach to achieve bounded-memory training and rendering of city-scale Gaussian splats at interactive rates.
[…]
Visual Abstract.
1. Introduction
Given a set of posed images of a 3D scene, the task of novel view synthesis (NVS) is to generate plausible images of the scene from unseen viewpoints. Early approaches achieved this via image-based blending (Zhang et al., 2019), but the introduction of Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) marked a breakthrough, enabling high-quality results by optimizing an implicit volumetric scene representation through a multi-layer perceptron. More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) extended this paradigm to an explicit representation: a set of Gaussian primitives that are efficiently rasterized using splatting techniques (Zwicker et al., 2001), replacing costly ray marching and allowing real-time rendering with fast convergence.
Despite these advances, both NeRF and 3DGS remain constrained by memory bottlenecks when applied to large-scale environments. Prior methods address this by dividing scenes into smaller chunks (Xu et al., 2023; Tancik et al., 2022; Li et al., 2024; Kerbl et al., 2024; Liu et al., 2024a; Chen et al., 2024b; Lin et al., 2024), training each independently before merging results. While chunking strategies mitigate memory usage during training, they introduces several key limitations:
  1. (1)
    View-chunk misalignment: Camera views often span multiple chunks, especially in open environments or multi-scale datasets (e.g., combining aerial and street-level images). This makes chunk boundaries arbitrary and hard to define, complicating scene partitioning and training.
  2. (2)
    Redundant overlap: To avoid artifacts at chunk boundaries, regions are typically trained with significant overlap, which increases memory usage and prolongs training time.
  3. (3)
    Asymmetric hardware demands: Although chunking reduces memory requirements during training, rendering may require all visible chunks in memory simultaneously—often exceeding the capacity of the original training setup.
The simplest and most robust alternative to chunking is to avoid splitting altogether. With A LoD of Gaussians, we introduce a seamless pipeline that enables training and rendering of ultra-large-scale scenes directly—on a single consumer-grade GPU—without any form of scene partitioning. To handle scenes that exceed available VRAM, we store all Gaussian data in CPU RAM and dynamically stream only those visible from the current training view into GPU memory. However, a single far-away view could still require access to the full scene. To address this, we construct a hierarchical Level-of-Detail (LoD) model inspired by Kerbl et al. (2024), loading detail proportional to view distance. Maintaining this hierarchy during training is non-trivial, as Gaussian properties evolve dynamically. We propose a novel hierarchy densification strategy, adapted from the MCMC-style spawning (Kheradmand et al., 2024), to support stable, progressive refinement. Efficient view-dependent selection from the hierarchy is challenging for large models. Instead of full tree traversal, we adopt Sequential Point Trees (SPTs) (Dachsbacher et al., 2003), originally developed for point cloud LoD rendering. Our Hierarichal SPT version allows us to compute the correct LoD cut efficiently for rendering individual views and camera paths. Finally, to reduce CPU-GPU data transfer overhead, we introduce a lightweight caching system that tracks recently used Gaussians and reuses them across training iterations. In summary, we make the following contributions:
  1. (1)
    We propose a novel hierarchy densification strategy that enables dynamic expansion and restructuring of Gaussian representations during training.
  2. (2)
    We leverage the Sequential Point Tree (SPT) data structure and adapt it for large-scale Gaussian Splatting to perform fast, parallelizable LoD cuts for efficient training and real-time rendering.
  3. (3)
    We demonstrate, for the first time, how external memory can be used to train and render ultra-large Gaussian scenes seamlessly on consumer-grade GPUs.
  4. (4)
    We design a caching and view scheduling system that exploits temporal coherence to minimize data transfer overhead and improve training throughput.
2. Related Work
Large Scale Reconstruction
Reconstructing large-scale scenes from images has long been a central challenge in visual computing. Traditional approaches relied on Structure-from-Motion pipelines, such as Building Rome in a Day (Agarwal et al., 2009) and COLMAP (Schönberger and Frahm, 2016), to recover geometry from unordered photo collections. Differentiable rendering techniques, notably NeRF (Mildenhall et al., 2020) and 3DGS (Kerbl et al., 2023), marked a paradigm shift by optimizing volumetric scene representations. Extensions of NeRF to large scenes typically employ scene partitioning (Tancik et al., 2022; Xu et al., 2023) or multi-GPU training strategies (Li et al., 2024). Similarly, most large-scale 3DGS pipelines adopt chunk-based training: Hierarchical 3DGS (Kerbl et al., 2024) trains chunks independently and then merges them into a global LoD hierarchy; CityGaussian (Liu et al., 2024a) combines chunked training with per-chunk LoD selection using LightGaussian (Fan et al., 2024); and VastGaussian (Lin et al., 2024) introduces decoupled appearance modeling and progressive partitioning. Horizon-GS (Jiang et al., 2024) integrates divide-and-conquer strategies with LoD mechanisms from (Ren et al., 2024), specifically targeting hybrid aerial/street-view datasets. Grendel (Zhao et al., 2024) avoids spatial chunking by distributing training images across GPUs, such that each device renders a disjoint screen region. Another research direction focuses on extracting geometric proxies from large-scale 3DGS scenes (Li et al., 2025; Liu et al., 2025; Chen et al., 2024b). These methods leverage TSDF fusion and geometric losses to generate multiple meshes, which are fused and rendered efficiently using traditional rasterization.
Compression
Lowering memory consumption is essential to scaling 3DGS to larger scenes. Compression techniques typically fall into three overlapping categories: compaction, which reduces the number of Gaussians; attribute compression, which reduces the storage per primitive; and structured representations, which organize data to improve compressibility. Pure compaction strategies include (Ren et al., 2024; Kim et al., 2024; Fang and Wang, 2024; Mallick et al., 2024; Cheng et al., 2024; Liu et al., 2024c); hybrid methods that also compress attributes include (Wang et al., 2024b; Papantonakis et al., 2024a; Fan et al., 2024; Girish et al., 2024); and pure attribute compression is addressed by (Niedermayr et al., 2024). Methods combining all three dimensions include (Chen et al., 2024a; Lee et al., 2024), while (Wang et al., 2024a; Ye et al., 2025; Navaneet et al., 2023; Liu et al., 2024b) emphasize attribute compression with structural priors. Finally, Lu et al. (2023) focus on hierarchical anchoring. These techniques are largely orthogonal to our work—many are applied post hoc or modify densification policies—and could be combined with our pipeline for further gains in memory and performance.
Level-of-Detail Rendering
Level-of-detail techniques reduce the complexity of distant scene content to accelerate rendering. In the context of 3DGS, LoD approaches have been explored for enabling efficient rendering on memory-constrained or mobile devices. Compression-based strategies include attribute quantization via codebooks, pruning of low-impact Gaussians, and adapting the degree of spherical harmonics per primitive (Papantonakis et al., 2024b; Niedermayr et al., 2024; Fan et al., 2024; Fang and Wang, 2024; Huang et al., 2025; Niemeyer et al., 2025; Seo et al., 2024). Scaffold-GS (Lu et al., 2023) introduced learned feature vectors anchored to reference Gaussians, with an MLP generating associated Gaussians at render time. This concept was extended to hierarchical LoD rendering in Octree-GS (Ren et al., 2024), enabling real-time control over detail levels through spatial subdivision.
3. Preliminaries
We briefly review 3D Gaussian Splatting and describe its memory usage characteristics. We then introduce the key data structures we used: Gaussian hierarchies and Sequential Point Trees.
3.1. 3D Gaussian Splatting
3D Gaussian Splatting (Kerbl et al., 2023) models a radiance field using a set of spatially distributed Gaussians, each with mean $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, RGB base colors $\mathbf{b}_{i}\in\mathbb{R}^{3}$ and covariance matrices $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$. The covariance is parameterized via a diagonal scaling matrix $\mathbf{S}_{i}=\text{diag}(s^{1}_{i},s^{2}_{i},s^{3}_{i})$ and an orthonormal rotation matrix $\mathbf{R}_{i}$:
$\boldsymbol{\Sigma}_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{\top}\mathbf{R}_{i}^{\top}.$
Each Gaussian also stores an opacity $\sigma_{i}$ and view-dependent color, modeled using spherical harmonics (SH) coefficients $\mathbf{f}_{i}^{d}$. The SH degree $d$ controls expressiveness, with each Gaussian requiring $\sum_{j=1}^{d}3\cdot(2j+1)$ parameters. For rendering, all $N$ Gaussians are sorted by distance to the camera and a discrete approximation of the volume rendering equation is evaluated for every pixel $\mathbf{x}$ with corresponding view direction $\mathbf{v}$:
$\mathbf{C}(\mathbf{x})=\sum_{i=1}^{N}\mathbf{c}_{i}(\mathbf{v})\alpha_{i}(\mathbf{x})\prod_{j=1}^{i-1}(1-\alpha_{j}(\mathbf{x})),$
where $\alpha_{j}$ is the opacity of the $j$-th Gaussian along the view ray:
$\alpha_{j}(\mathbf{x})=\sigma_{j}e^{-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu^{\prime}}_{j})\boldsymbol{\Sigma^{\prime}}(\mathbf{x}-\boldsymbol{\mu^{\prime}}_{j})^{T}}.$
Here $\boldsymbol{\mu^{\prime}}$ and $\boldsymbol{\Sigma^{\prime}}$ denote the projected 2D mean and covariance on the image plane, obtained by applying an affine approximation of the projective transform (Zwicker et al., 2001). The rendering pipeline is fully differentiable, enabling end-to-end optimization with ADAM using a combination of $\mathcal{L}_{1}$ and SSIM losses.
3.2. Memory in 3DGS
Standard 3DGS pipelines store the full set of per-Gaussian attributes, training images, and optimizer state in GPU memory (VRAM). Figure 15 summarises the amount of memory that the various components of a single Gaussian require. Additional temporary allocations occur during forward and backward passes (e.g., for sorting and gradient accumulation). This overhead varies with the scale of the Gaussians and effectiveness of culling strategies, but can be roughly upper-bounded by around 800 bytes per Gaussian in practice. This limits training to roughly $500$K Gaussians per GB of VRAM, imposing strict constraints on the detail and extent of reconstructions.
3.3. Gaussian Hierarchies and Sequential Point Trees
Gaussian hierarchies, introduced in (Kerbl et al., 2024), recursively merge nearby Gaussians into a tree, where each non-leaf node approximates its children, and leaves correspond to the original Gaussians. A cut is defined by a condition $c_{\text{hier}}(i,\text{cam})$ evaluated in a breadth-first search (BFS). If a node satisfies the cut condition, it is added to the cut set and its children are skipped; otherwise, the BFS continues. A proper cut set includes no parents or children of any included node and thus provides a view-adaptive LoD representation.
The cut condition used in (Kerbl et al., 2024) is:
$c_{\text{hier}}(i,\text{cam})=\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i),\quad m_{d}(i)=\frac{T}{\max_{j}s_{i}^{j}},$
where $\mathbf{p}_{cam}$ is the camera position, $T$ is a global LoD threshold, and $m_{d}(i)$ is the minimum acceptable distance for viewing Gaussian $i$. This ensures that if $i$ is in the cut set, its parent failed the condition:
$m_{d}(\text{parent}(i))>\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$
Sequential Point Trees (SPTs) (Dachsbacher et al., 2003), originally developed for point cloud LoD rendering, can be adapted for Gaussians. They enforce a more constrained cut condition:
$c_{\text{SPT}}(i,\text{cam})=m_{d}(\text{parent}(i))>\left\|\boldsymbol{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$
This condition is evaluated for all Gaussians in parallel, using the shared root-camera distance $\left\|\boldsymbol{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$. It requires storing only sorted pairs $\big{(}m_{d}(i),m_{d}(\text{parent}(i))\big{)}$, significantly reducing memory compared to full Gaussian hierarchies. Note that cuts are guaranteed to be proper, with nodes where $m_{d}(i)<m_{d}(\text{parent}(i))$ never being selected for the cut. To optimize cuts, Gaussians are sorted by $m_{d}(\text{parent}(i))$ in descending order. A binary search determines the cutoff index $N$, above which Gaussians are too fine to be rendered. The downside of the approach is that all Gaussians in the same SPT share the same level of detail dictated by the distance from the camera to the root node. This can lead to Gaussians with $m_{d}(i)>\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}$ being rendered, even though they would be too coarse for the current view. To counteract this issue, we define:
$M_{d}(i)=m_{d}(i)+\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2},$
as a conservative minimum distance function. By the triangle inequality, selecting Gaussians satisfying $M_{d}(i)\leq\left\|\boldsymbol{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$ guarantees $m_{d}(i)\leq\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}$. In turn, this means that Gaussians that are further away from the camera than the root node will be selected at a higher level of detail. SPTs are best suited for tightly grouped Gaussians observed from distances greater than their mutual spacing. Their compact memory footprint and parallel evaluation make them well-suited for large-scale scenes. Figure 2 visualizes both hierarchy types and their LoD cuts for a toy example.
Figure 2. A Sequential Point Tree and Gaussian hierarchy represent the same 5 Gaussians at varying levels of detail shown with four possible hierarchy cuts in red. Horizontal lines in the SPT show the binary search result and horizontal lines the distance cut.
[…]
[SPT vs Gaussian hierarchy]Sequential Point Tree and Gaussian hierarchy datastructure for a set of 9 Gaussians along with various levels of detail and their cuts
4. Method
An overview of the training and densification process can be found in Figure 7, where we enumerate the individual steps of our training and densification process: To train models exceeding GPU memory limits, all Gaussian attributes are stored in CPU RAM and streamed to the GPU on demand for each training view. To accelerate the hierarchy cut, we store a copy of only the tree structure in VRAM, where larger subtrees have been replaced by SPTs, forming a hierarchical SPT (HSPT). To minimize costly transfers between RAM and VRAM, we keep track of which SPTs are currently loaded and at which detail, and cache them in GPU memory. Only if an SPT is not present in this GPU cache at a similar level of detail, will it be loaded from RAM. Densification occurs on the CPU by adding new leaf nodes to the hierarchy and respawning low-opacity leaf nodes. The densified hierarchy is then converted back to an HSPT and transferred to the GPU for a new round of training iterations.
4.1. Initialization
Following Kerbl et al. (2024), we initialize the Gaussian model from a sparse point cloud, augmented with skybox points. This initial representation is small enough to fully reside in GPU memory and is trained for 100 000100000100\,000100 000 iterations with densification disabled. The goal of this phase is to establish a stable global scene structure before hierarchy construction. After this initial optimization, we build a binary Gaussian hierarchy, with the trained Gaussians as leaf nodes and parent nodes representing merged approximations of their children.
4.2. Gradient Propagation
To update the properties of Gaussians, including base color, SH coefficients, position, and covariance, we follow the standard 3DGS error propagation (Kerbl et al., 2023). However, we operate only on the cut set chosen for each training view. Thus, gradients may, in general, be propagated to Gaussians in the middle of the hierarchy.
4.3. Densification
Densifying a Level-of-Detail representation presents a unique challenge: the underlying hierarchical structure must evolve continuously during training. Prior works circumvent this issue by constructing LoD hierarchies only after chunk-level training and densification are complete.
We take inspiration from 3DGS-MCMC (Kheradmand et al., 2024), which selects Gaussians probabilistically (weighted by opacity) to be ‘split’, replacing it by two new Gaussians which together should appear similarly to the original split Gaussian. Noteably, this resembles the way a parent node in a Gaussian hierarchy approximates its children. Therefore, we adopt this approach and ‘spawn’ two new child nodes for a leaf instead of splitting a Gaussian, increasing the size of the hierarchy with minimal artifacts.
Instead of pruning, Kheradmand et al. (2024) declares Gaussians below a certain opacity threshold as ‘dead’, and respawns them at the position of a high-opacity Gaussian. We propose a similar strategy: when a leaf node dies, its parent is replaced by its sibling node; the dead leaf node and its parent are then respawned as children to another node, which is selected to be densified. See Figure 3 for an overview of the two hierarchy densification operations. Together, they ensure that the hierarchy can be expanded during training in a stable manner and rebalanced as required. Avoiding destructive resets, as advocated by Kerbl et al. (2023), is particularly important in our setting, as many Gaussians will no longer conform to the hierarchical structure after opacity is restored to normal.
Figure 3. Example of densifying and respawning leaf nodes.
4.4. The hierarchical SPT datastructure
BFS
Computing the cut set of a large Gaussian hierarchy is a costly operation that must be performed for every frame. A straightforward solution is to run a breadth-first search (BFS) from the root. This guarantees a proper cut and enables early pruning of large subtrees, e.g., via frustum culling. However, graph traversal is not well suited to parallel execution on the GPU.
Parallel Cut
To enable GPU-accelerated cuts, Kerbl et al. (2024) evaluate the cut condition in parallel for each Gaussian:
$\Big{(}m_{d}(i)<\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}\Big{)}\land\Big{(}m_{d}(parent(i))\geq\left\|\boldsymbol{\mu}_{\text{parent}(i)}-\mathbf{p}_{cam}\right\|_{2}\Big{)},$
where any Gaussian that is sufficiently small at its current camera distance and whose parent is too large to be rendered, should be part of the cut set. This produces a proper cut under the assumption that child Gaussians always have a smaller minimal distance than their parents:
$\forall i:m_{d}(i)<m_{d}(parent(i)).$
This is generally valid when hierarchies are constructed after training, since parent Gaussians represent coarser approximations of their children. However, when the hierarchy is modified during training and densification, optimization can cause this assumption to break—leading to invalid cut sets and degenerate hierarchies that worsen over time.
Hierarchical SPT (HSPT)
We present the hierarchical SPT data structure, which combines the benefits of both approaches. To construct it, we cut it using a BFS on the condition
$c_{HSPT}(i)=s_{i}^{1}\cdot s_{i}^{2}\cdot s_{i}^{3}<\texttt{size}$
with volume threshold size. The resulting cut set $\mathbb{C}_{\text{HSPT}}$ partitions the hierarchy into the upper hierarchy, which includes all Gaussians with volume greater than size, and the lower hierarchy, consisting of the subtrees rooted at the nodes in the cut set.
The volume of each root node in the lower hierarchy is now bounded by size, which also roughly bounds the extent of all Gaussians in the subtree. This provides an upper bound on the error introduced if the subtree is converted into an SPT. Consequently, each subtree of sufficient size in $\mathbb{C}_{\text{HSPT}}$ can be transformed into a Sequential Point Tree to accelerate cut computation. The HSPT-based cutting process then proceeds in two steps: first, a BFS on the upper hierarchy selects the required nodes and leaf/SPT subtrees for the current view. Second, each selected SPT is cut according to the camera’s distance to its root node. Together, these yield the full set of Gaussians needed for rendering the current frame.
The construction and cutting process of an HSPT is illustrated in Figure 4. The SPTs used to render a MatrixCity-Scale frame are shown in Figure 11, and Figure 10 demonstrates the smooth LoD transitions enabled by the hierarchical structure.
Figure 4. A Gaussian hierarchy is converted to an HSPT by cutting according to Gaussian volume and converting sufficiently large subtrees to SPTs. The HSPT can then be cut in a 2-step process.
Rebuilding the hierarchical SPT for every training iteration would eliminate any performance benefit. Instead, we exploit the fact that the minimum distance $m_{d}$ evolves slowly during optimization and thus only needs to be updated infrequently. In practice, we rebuild the HSPT only after each densification step.
This infrequent recomputation allows us to use a more accurate—albeit more expensive—minimum distance metric than the inverse of maximal scale. Specifically, we define:
$m_{d}^{\prime}(i)=\frac{T}{\sqrt{s_{i}^{1}\cdot s_{i}^{2}+s_{i}^{1}\cdot s_{i}^{3}+s_{i}^{2}\cdot s_{i}^{3}}},$
which corresponds to the inverse square root of the surface area of the Gaussian ellipsoid (up to a constant factor). This better captures the perceived size of anisotropic Gaussians, especially those that are significantly elongated in one or more directions.
Frustum Culling
The major benefits of using BFS to cut the upper hierarchy are the guarantee of a proper cut and early culling of subtrees. To this end, we frustum cull every node considered in the BFS by checking if a sphere around the Gaussians with radius equal to $(3\cdot\max_{j}s_{i}^{j})$ intersects with the view frustum. Using the Gaussian scale as a proxy for the extent of its entire subtree is not perfectly accurate, but comparisons to using a full bounding sphere hierarchy showed no discernable difference in practice. It should be noted that Gaussians are implicitly frustum culled during rasterization, but this early culling accelerates the cutting procedure and significantly decreases the number of Gaussians that need to be loaded from RAM, as can be seen in Figure 9.
4.5. Caching on the GPU
Loading Gaussian data from RAM is a costly operation that can become a significant bottleneck during large-scale training. To mitigate this, we maintain a GPU-resident cache of Gaussians that are likely to be reused across consecutive training views. However, checking the cache for every individual Gaussian would introduce nontrivial overhead. Once again, SPTs offer an efficient alternative.
Rather than caching individual Gaussians, we store the Gaussians from SPT cuts along with the cached distance from the camera to the root of each SPT, denoted $\bar{d}^{j}$ for the $j$th SPT. During rendering, when the upper hierarchy is cut and the required SPTs identified, we compute $d^{j}=\|\boldsymbol{\mu}_{\text{root}(j)}-\mathbf{p}_{\text{cam}}\|_{2}$ and check whether a matching cut is already cached. This check is performed using a simple distance ratio tolerance:
$D_{\text{min}}\leq\frac{d^{j}}{\bar{d}^{j}}\leq D_{\text{max}}.$
Here, $D_{\text{max}}$ defines the allowable range for using coarser-than-optimal detail, while $D_{\text{min}}$ limits how much finer detail can be tolerated. If this condition is met, the cached SPT cut is reused, avoiding a costly RAM-to-GPU transfer. While this heuristic introduces slight variability in rendered detail—since the LoD may depend on the cache state—we find that this stochasticity actually improves training robustness. In particular, the subtle variation in detail across views discourages overfitting to fixed camera distances and promotes better generalization across scales.
For each training view, visible Gaussians are assembled from the upper hierarchy, the cached SPTs, and the skybox Gaussians (which remains in VRAM). Uncached SPTs are streamed from RAM. After each training iteration, newly loaded SPTs are added to the cache.
To bound VRAM usage, the we use a least-recently-used (LRU) write-back policy. When a memory threshold is exceeded, entries are written back to RAM. Additionally, to prevent overfitting to persistent cache entries, the entire cache is flushed every 1 00010001\,0001 000 iterations. Figure 5 illustrates the caching process across two frames.
Figure 5. Caching strategy overview. Gaussians required for the current training view are assembled from three sources: the upper tree, newly loaded SPT cuts from RAM, and cache hits. After optimization, newly accessed SPTs are added to the GPU cache.
View Selection
In large-scale scenes, the GPU cache typically covers only a small fraction of the overall geometry, leading to sparse cache hits. To improve cache utilization, we prioritize spatial locality by selecting successive training views that are geographically close to the current one, thereby maximizing Gaussian reuse.
To this end, we precompute a $k$-nearest-neighbor graph over all training view positions, where edge weights $w_{ij}$ correspond to the Euclidean distance between views $i$ and $j$. The next training view $j$ is then sampled from the $k$-nearest neighbors of the current view $i$ according to the distribution:
$\mathbb{P}(j\mid i)\propto\frac{1}{w_{ij}+W},$
where $W$ is a constant, controlling the degree of exploration.
However, care must be taken when modifying the order of training views, as deviating from uniform random sampling may introduce bias. Thus, we inject a random view every 12 iterations, which we find sufficient to preserve generalization performance.
4.6. Memory Layout
Figure 6 illustrates the peak memory usage for a single training iteration of a 60-million-Gaussian hierarchy on the MatrixCity-Scale dataset. The output frame from this iteration is shown in Figure 12. The majority of RAM usage is consumed by per-Gaussian properties and their corresponding ADAM optimizer states. In contrast, the hierarchy structure itself accounts for less than 10% of the total RAM footprint. On the GPU, the SPT metadata for all 60 million Gaussians occupies just 680 MB of VRAM and the upper hierarchy negligible 24 MB.
Even in wide-angle aerial views, only a subset of the scene is actively loaded into GPU memory. In the example shown, 2.2 million Gaussians are rendered directly, while an additional 2.4 million are retained in the cache for future use. The bulk of GPU memory is instead consumed by temporary allocations for rasterization and optimization, which scale with the number of Gaussians rendered. Therefore, minimizing the number of active Gaussians is critical for staying within the VRAM budget.
The remaining GPU memory usage consists of auxiliary data, including cache management, hierarchy cut tracking, training and ground-truth images, as well as PyTorch overhead. Figure 15 provides a breakdown of memory usage for Gaussian properties and hierarchy data, and how these are distributed across CPU and GPU.
Figure 6. Peak memory consumption of CPU and GPU for a training iteration on BigCity-Scale with 60 million Gaussians.
5. Evaluation
Choice of dataset
Our method is designed to enable seamless training and rendering on large-scale scenes comprising thousands of views captured at vastly different scales. Unfortunately, most existing datasets of sufficient size contain either street-level or aerial views—but not both. Additionally, at this scale and density, current Structure-from-Motion (SfM) pipelines often fail to reliably recover camera poses and sparse point clouds.
To overcome these challenges, we use the synthetic MatrixCity dataset (Li et al., 2023), which provides dense multi-scale views along with ground-truth camera poses and sparse geometry. For our main benchmark, MatrixCity-Scale (MC-Scale), we aggregate 11 466 street-view images, 3 208 aerial views, and 533 additional high-altitude views we generate manually. We evaluate reconstruction quality on a separate set of 1 694 test images. The diversity in image scale is visualized in Figure 8. To validate our method on real-world data, we also include results on the largest public dataset from Kerbl et al. (2024), the Campus scene.
Choice of Comparison
Selecting appropriate baselines is challenging. Standard 3DGS variants perform poorly at large scales, while most divide-and-conquer approaches—such as Hierarchical 3DGS, CityGS, and HorizonGS—assume access to 80 GB GPUs. Despite access to multiple clusters, we were limited to 40 GB A100s, highlighting the hardware demands of these methods.
To our knowledge, only Hierarchical 3DGS has trained successfully on the Campus dataset, and no prior work has demonstrated large-scale training on MatrixCity with combined aerial and street views. Due to resource constraints, we were unable to train CityGS on either dataset and adopt Hierarchical 3DGS as our main baseline.
For additional reference, we evaluate 3DGS-MCMC, modified to support large-scale training by limiting densification, streaming images from disk, and reducing the SH degree. Most chunk-based methods, including Kerbl et al. (2024), require feature correspondences, which are not provided in MatrixCity. To enable comparison, we reconstructed a sparse point cloud using COLMAP and the provided camera poses. This less stable point cloud (visualized in Figure 8) was used for initialization.
On the Campus scene, we report results for the official test split (Hierarchical single) and the full pretrained model. Consequently, we evaluate reconstruction quality on every 100th training view. We train a smaller model (38M) and a larger (80M) model. Rendering Hierarchical 3DGS required reducing SH degree to 1—even on 40 GB VRAM; ours also uses SH degree 1. Given the limitations of numerical metrics in these settings, we strongly encourage visual comparison via the supplementary videos, which in our opinion favour Hierarchical 3DGS. Except for the results of Hierarchical 3DGS, all training and rendering was done using an RTX 3090 GPU.
5.1. Results
Quantitative comparisons are shown in Table 1. A LoD of Gaussians successfully reconstructs the full MC-Scale scene, where other methods struggle due to the scene’s size and diverse camera viewpoints. Notably, Kerbl et al. (2024) reconstructs individual chunks accurately, but the merging process introduces ”floaters”—isolated Gaussians not associated with meaningful geometry—which obscure most test views (see Figure 13 and 14). Similarly, 3DGS-MCMC suffers from aggressive opacity and scale-based pruning, causing a large fraction of Gaussians to vanish entirely.
On the Campus dataset, Hierarchical 3DGS performs well, as its chunk-based approach aligns with the dataset’s single, street-level trajectory. Their single-chunk evaluation shows strong metrics. However, when evaluating their full model with reduced SH degree (to fit within 40 GB VRAM), we observe a clear drop in numerical quality. That said, the visual results remain competitive—see the supplemental video. In our case, both low- and high-resolution variants achieve similar quantitative scores despite visible differences, due to Hierarchical 3DGS evaluating on training views.
For context, Hierarchical 3DGS splits the MC-Scale scene into 94 chunks, each trained for 45 000 iterations, totaling over 4.23 million steps and nearly four full days of compute time on our machine. By contrast, our approach trains the full scene end-to-end in just 250 000 iterations, completing in approximately 15 hours.
Memory
A major limitation of chunk-based methods is the memory overhead during rendering. Although it also implements a form of GPU streaming, we were unable to run the viewer provided by Kerbl et al. (2024) on either dataset using the 40 GB VRAM A100, which is consistent with the limitations they report. In contrast, our method renders interactive flythroughs of MatrixCityScale and Campus (38M), see supplemental, while only requiring 8 GB of GPU memory (and 16 GB for Campus (80M)).
Table 1. Full Novel View Synthesis results. * use of the suboptimal COLMAP version for MC-Scale; adjusted SH degree to match ours and reduce memory to enable rendering, use of higher SH degree and only single chunk, reproduced from (Kerbl et al., 2024).
MC-Scale (15k images) Campus (22k images)
[…]
Ours* 20.63 0.668 0.282 40M 22.83 0.713 0.248 38M
Hierarchical*,‡ 13.77 0.519 0.559 82M 17.76 0.601 0.424 80M
Ours 21.73 0.712 0.213 60M 22.86 0.725 0.237 80M
Hierarchical single 24.61 0.807 0.331 -
3DGS-MCMC 13.16 0.4481 0.586 6M 15.11 0.600 0.580 6M
Table 2. Ablations. Average frame times for rendering camera paths and average iteration times during training with and without caching Gaussians. Then, average times to cut the hierarchy during rendering with the normal BFS approach and HSPT. 80M Campus was trained longer than 38M.
MC-Scale Campus
38M 80M
Render Full 48.1 ms 47.1 ms 83.2 ms
Render w/o Cache 119.4 ms 92.6 ms 222.3ms
Render w/o Frustum Culling 52.3 ms 38.3 ms 110.2 ms
Train Full 156 ms 205 ms
Train w/o Cache 471 ms 244 ms
Train w/o Frustum Culling 685 ms 312 ms
HSPT Cut (Render) 31.9 ms 31.3 ms 36.5 ms
BFS Cut (Render) 47.8 ms 40.0 ms 53.7 ms
5.2. Ablations
We assess the contribution of key components through an ablation study, using recorded camera paths across all scenes (see supplemental video). Table 2 reports average frame times over these paths. Caching significantly improves performance, roughly doubling the framerate across scenes. Frustum culling is particularly beneficial for large-scale models but may incur overhead in smaller scenes. To evaluate cut efficiency, we compare the time required to compute the visible set using either full hierarchy BFS or our HSPT-based approach. HSPT consistently yields faster cut times. For training, we measure average iteration durations over 1 000 steps. Here, frustum culling again proves essential, substantially reducing the number of Gaussians loaded and processed per view.
View Selection
On the MatrixCity-Scale dataset (250K iterations, up to 60M Gaussians), enabling view selection reduces the number of Gaussians loaded from RAM per frame from 355K to 219K—a $35\%$ reduction in memory load. Reconstruction quality is unaffected, with PSNR and SSIM improving slightly by 0.04 and 0.05, respectively.
6. Discussion and Outlook
A LoD of Gaussians enables seamless training and rendering of ultra-large 3DGS models on consumer hardware. By storing Gaussian data in external memory and streaming it on demand, our method avoids the pitfalls of chunk-based pipelines. The hierarchical SPT structure accelerates LoD selection and remains robust to ongoing training changes. Combined with caching and view selection, our system significantly reduces out-of-core overhead. Together, these components enable efficient reconstruction and rendering at scale, demonstrated on the MatrixCity-Scale dataset.
Limitations and Future Work
A primary limitation of our method lies in initialization. Accurate estimation of camera poses and sparse point clouds remains challenging at large scale, particularly for real-world datasets with sparse or inconsistent coverage. Future work could explore improved initialization strategies or joint optimization of poses and Gaussians during early training.
Our system also requires roughly 1 GB of RAM per million Gaussians, which—while more efficient than prior methods—still constrains scalability. Loading from disk is feasible in our experiments, though at the cost of a 10×10$\times$10 × slowdown. Preliminary exploration has also shown that given access to sufficient RAM and training time, our method could scale far beyond what is presented here. A 150 million Gaussian model has been trained successfully for MatrixCity-Scale on 24 GB of VRAM, though the improvement to image quality was minor. We suspect that even more training views are required to properly leverage such a vast number of Gaussians.
Rendering could be further optimized by avoiding per-frame hierarchy cut recomputation and enabling asynchronous streaming. While frustum culling reduces memory load in most views, it is ineffective when the entire scene falls inside the frustum. Occlusion culling could address this by skipping entire SPTs before loading.
Overall, we believe that out-of-core 3D Gaussian Splatting is a promising direction for scaling radiance field methods to city-scale and beyond, without the need for specialized hardware.

### v3

@@
 A LoD of Gaussians: Unified Training and Rendering for Ultra-Large Scale Reconstruction with External Memory
-Abstract.
+Abstract
 Gaussian Splatting has emerged as a high-performance technique for novel view synthesis, enabling real-time rendering and high-quality reconstruction of small scenes.
@@
 Moreover, rendering remains fundamentally limited by GPU memory, as all visible chunks must reside in VRAM simultaneously.
-We introduce A LoD of Gaussians, a framework for training and rendering ultra-large-scale Gaussian scenes on a single consumer-grade GPU—without partitioning.
+We introduce A LoD of Gaussians, a framework for training and rendering ultra-large-scale Gaussian scenes on a single consumer-grade GPU without partitioning.
 Our method stores the full scene out-of-core (e.g., in CPU memory) and trains a Level-of-Detail (LoD) representation directly, dynamically streaming only the relevant Gaussians.
-A hybrid data structure combining Gaussian hierarchies with Sequential Point Trees enables efficient, view-dependent LoD selection, while a lightweight caching and view scheduling system exploits temporal coherence to support real-time streaming and rendering.
+A hybrid data structure combining Gaussian hierarchies with Sequential Point Trees enables efficient, view-dependent LoD selection, while a lightweight caching and view scheduling system exploits temporal coherence to minimize the loading overhead.
 Together, these innovations enable seamless multi-scale reconstruction and interactive visualization of complex scenes—from broad aerial views to fine-grained ground-level details.
-Level of Detail, Gaussian Splatting, Large-Scale Reconstruction
-isbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Rasterizationccs: Computing methodologies Visibilityccs: Computing methodologies Machine learning
-Figure 1.
-We introduce a fully hierarchical 3D Gaussian representation trained directly across unstructured, multi-scale image sets—including street-level and aerial views—without scene partitioning.
-Our method maintains a consistent global scene model, eliminating boundary artifacts typical of chunked approaches.
-A hybrid Level-of-Detail system combines Gaussian hierarchies with Sequential Point Trees, enabling dynamic, view-dependent streaming and LoD selection.
-The entire model resides in external memory, with only a small, adaptive subset loaded on demand, allowing real-time rendering and training of 60M+ Gaussians on consumer GPUs ($\leq 24GB$ VRAM).
-This is the first approach to achieve bounded-memory training and rendering of city-scale Gaussian splats at interactive rates.
-[…]
-Visual Abstract.
-1. Introduction
+1 Introduction
 Given a set of posed images of a 3D scene, the task of novel view synthesis (NVS) is to generate plausible images of the scene from unseen viewpoints.
-Early approaches achieved this via image-based blending (Zhang et al., 2019), but the introduction of Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) marked a breakthrough, enabling high-quality results by optimizing an implicit volumetric scene representation through a multi-layer perceptron.
-More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) extended this paradigm to an explicit representation: a set of Gaussian primitives that are efficiently rasterized using splatting techniques (Zwicker et al., 2001), replacing costly ray marching and allowing real-time rendering with fast convergence.
+Early approaches achieved this via image-based blending (DeepBlending), but the introduction of Neural Radiance Fields (NeRF) (NeRF) marked a breakthrough, enabling high-quality results by optimizing an implicit volumetric scene representation through a multi-layer perceptron.
+More recently, 3D Gaussian Splatting (3DGS) (3DGS) extended this paradigm to an explicit representation: a set of Gaussian primitives that are efficiently rasterized using splatting techniques (zwicker2001EWA), replacing costly ray marching and allowing real-time rendering with fast convergence.
 Despite these advances, both NeRF and 3DGS remain constrained by memory bottlenecks when applied to large-scale environments.
-Prior methods address this by dividing scenes into smaller chunks (Xu et al., 2023; Tancik et al., 2022; Li et al., 2024; Kerbl et al., 2024; Liu et al., 2024a; Chen et al., 2024b; Lin et al., 2024), training each independently before merging results.
-While chunking strategies mitigate memory usage during training, they introduces several key limitations:
-  1. (1)
+Prior methods address this by dividing scenes into smaller chunks (GridNeRF; BlockNeRF; Hierarchical; CityGaussian; GigaGS; VastGaussian), training each independently before merging results.
+While chunking strategies mitigate memory usage during training, they introduce several key limitations:
+  1. 1.
     View-chunk misalignment: Camera views often span multiple chunks, especially in open environments or multi-scale datasets (e.g., combining aerial and street-level images).
 This makes chunk boundaries arbitrary and hard to define, complicating scene partitioning and training.
-  2. (2)
+We identify two artifacts caused by view-chunk misalignment: Chunk bleeding occurs when Gaussians extend out of their current chunk and obscure neighbouring chunks after merging.
+When an occluder on a training image is not part of the current chunk, its ”ghostly” outline will be trained into the wrong chunk, leading to Chunk ghosting.
+Examples of these artifacts can be seen in Figures6 and 11.
+  2. 2.
     Redundant overlap: To avoid artifacts at chunk boundaries, regions are typically trained with significant overlap, which increases memory usage and prolongs training time.
-  3. (3)
+  3. 3.
     Asymmetric hardware demands: Although chunking reduces memory requirements during training, rendering may require all visible chunks in memory simultaneously—often exceeding the capacity of the original training setup.
 The simplest and most robust alternative to chunking is to avoid splitting altogether.
-With A LoD of Gaussians, we introduce a seamless pipeline that enables training and rendering of ultra-large-scale scenes directly—on a single consumer-grade GPU—without any form of scene partitioning.
+With A LoD of Gaussians, we introduce a seamless pipeline that enables training and rendering of ultra-large-scale scenes directly on a single consumer-grade GPU, without any form of scene partitioning (see Figure 7).
 To handle scenes that exceed available VRAM, we store all Gaussian data in CPU RAM and dynamically stream only those visible from the current training view into GPU memory.
-However, a single far-away view could still require access to the full scene.
-To address this, we construct a hierarchical Level-of-Detail (LoD) model inspired by Kerbl et al.
-(2024), loading detail proportional to view distance.
+Still, a single distant view could require access to the full scene.
+To address this, we construct a hierarchical Level-of-Detail (LoD) model inspired by Hierarchical, loading detail proportional to view distance.
 Maintaining this hierarchy during training is non-trivial, as Gaussian properties evolve dynamically.
-We propose a novel hierarchy densification strategy, adapted from the MCMC-style spawning (Kheradmand et al., 2024), to support stable, progressive refinement.
+We propose a novel hierarchy densification strategy, adapted from the MCMC-style spawning (3DGSasMCMC), to support stable, progressive refinement.
 Efficient view-dependent selection from the hierarchy is challenging for large models.
-Instead of full tree traversal, we adopt Sequential Point Trees (SPTs) (Dachsbacher et al., 2003), originally developed for point cloud LoD rendering.
-Our Hierarichal SPT version allows us to compute the correct LoD cut efficiently for rendering individual views and camera paths.
+Instead of full tree traversal, we adopt Sequential Point Trees (SPTs) (SPT), originally developed for point cloud LoD rendering.
+Our Hierarchical SPT version allows us to compute the correct LoD cut efficiently for rendering individual views and camera paths.
 Finally, to reduce CPU-GPU data transfer overhead, we introduce a lightweight caching system that tracks recently used Gaussians and reuses them across training iterations.
 In summary, we make the following contributions:
-  1. (1)
-    We propose a novel hierarchy densification strategy that enables dynamic expansion and restructuring of Gaussian representations during training.
-  2. (2)
-    We leverage the Sequential Point Tree (SPT) data structure and adapt it for large-scale Gaussian Splatting to perform fast, parallelizable LoD cuts for efficient training and real-time rendering.
-  3. (3)
+  1. 1.
+    We propose a novel hierarchy densification strategy that enables dynamic expansion and restructuring of Gaussian LoD representations during training.
+  2. 2.
+    We leverage the Sequential Point Tree (SPT) data structure and adapt it for large-scale Gaussian Splatting to perform fast, parallelizable LoD cuts for efficient training and rendering.
+  3. 3.
     We demonstrate, for the first time, how external memory can be used to train and render ultra-large Gaussian scenes seamlessly on consumer-grade GPUs.
-  4. (4)
+  4. 4.
     We design a caching and view scheduling system that exploits temporal coherence to minimize data transfer overhead and improve training throughput.
-2. Related Work
+  5. 5.
+    We verify trough extensive evaluations that the pipeline succeeds in reconstructing huge datasets with complex camera distribution, where SOTA divide-and-conquer methods fail even with larger VRAM budgets.
+2 Related Work
 Large Scale Reconstruction
 Reconstructing large-scale scenes from images has long been a central challenge in visual computing.
-Traditional approaches relied on Structure-from-Motion pipelines, such as Building Rome in a Day (Agarwal et al., 2009) and COLMAP (Schönberger and Frahm, 2016), to recover geometry from unordered photo collections.
-Differentiable rendering techniques, notably NeRF (Mildenhall et al., 2020) and 3DGS (Kerbl et al., 2023), marked a paradigm shift by optimizing volumetric scene representations.
-Extensions of NeRF to large scenes typically employ scene partitioning (Tancik et al., 2022; Xu et al., 2023) or multi-GPU training strategies (Li et al., 2024).
-Similarly, most large-scale 3DGS pipelines adopt chunk-based training: Hierarchical 3DGS (Kerbl et al., 2024) trains chunks independently and then merges them into a global LoD hierarchy; CityGaussian (Liu et al., 2024a) combines chunked training with per-chunk LoD selection using LightGaussian (Fan et al., 2024); and VastGaussian (Lin et al., 2024) introduces decoupled appearance modeling and progressive partitioning.
-Horizon-GS (Jiang et al., 2024) integrates divide-and-conquer strategies with LoD mechanisms from (Ren et al., 2024), specifically targeting hybrid aerial/street-view datasets.
-Grendel (Zhao et al., 2024) avoids spatial chunking by distributing training images across GPUs, such that each device renders a disjoint screen region.
-Another research direction focuses on extracting geometric proxies from large-scale 3DGS scenes (Li et al., 2025; Liu et al., 2025; Chen et al., 2024b).
+Traditional approaches relied on Structure-from-Motion (SfM) pipelines to recover geometry from unordered photo collections (RomeInADay; COLMAP).
+Differentiable rendering techniques, notably NeRF (NeRF) and 3DGS (3DGS), marked a paradigm shift by optimizing volumetric scene representations.
+Extensions of NeRF to large scenes typically employ scene partitioning (BlockNeRF; GridNeRF) or multi-GPU training strategies (NerfXL).
+Similarly, most large-scale 3DGS pipelines adopt chunk-based training: Hierarchical-3DGS (H-3DGS) (Hierarchical) trains chunks independently and then merges them into a global LoD hierarchy; CityGaussian (CityGaussian) combines chunked training with per-chunk LoD selection using LightGaussian (LightGaussian); and VastGaussian (VastGaussian) introduces decoupled appearance modeling and progressive partitioning.
+Horizon-GS (HorizonGS) integrates divide-and-conquer strategies with LoD mechanisms from OctreeGS, specifically targeting hybrid aerial/street-view datasets.
+GrendelGS (Grendel) avoids spatial chunking by distributing training images across GPUs, such that each device renders a disjoint screen region.
+Another research direction focuses on extracting geometric proxies from large-scale 3DGS scenes (ULSRGS; CityGaussianV2; GigaGS).
 These methods leverage TSDF fusion and geometric losses to generate multiple meshes, which are fused and rendered efficiently using traditional rasterization.
-Compression
-Lowering memory consumption is essential to scaling 3DGS to larger scenes.
-Compression techniques typically fall into three overlapping categories: compaction, which reduces the number of Gaussians; attribute compression, which reduces the storage per primitive; and structured representations, which organize data to improve compressibility.
-Pure compaction strategies include (Ren et al., 2024; Kim et al., 2024; Fang and Wang, 2024; Mallick et al., 2024; Cheng et al., 2024; Liu et al., 2024c); hybrid methods that also compress attributes include (Wang et al., 2024b; Papantonakis et al., 2024a; Fan et al., 2024; Girish et al., 2024); and pure attribute compression is addressed by (Niedermayr et al., 2024).
-Methods combining all three dimensions include (Chen et al., 2024a; Lee et al., 2024), while (Wang et al., 2024a; Ye et al., 2025; Navaneet et al., 2023; Liu et al., 2024b) emphasize attribute compression with structural priors.
-Finally, Lu et al.
-(2023) focus on hierarchical anchoring.
-These techniques are largely orthogonal to our work—many are applied post hoc or modify densification policies—and could be combined with our pipeline for further gains in memory and performance.
 Level-of-Detail Rendering
-Level-of-detail techniques reduce the complexity of distant scene content to accelerate rendering.
+Level-of-detail techniques reduce the geometric complexity of distant scene content to accelerate rendering.
 In the context of 3DGS, LoD approaches have been explored for enabling efficient rendering on memory-constrained or mobile devices.
-Compression-based strategies include attribute quantization via codebooks, pruning of low-impact Gaussians, and adapting the degree of spherical harmonics per primitive (Papantonakis et al., 2024b; Niedermayr et al., 2024; Fan et al., 2024; Fang and Wang, 2024; Huang et al., 2025; Niemeyer et al., 2025; Seo et al., 2024).
-Scaffold-GS (Lu et al., 2023) introduced learned feature vectors anchored to reference Gaussians, with an MLP generating associated Gaussians at render time.
-This concept was extended to hierarchical LoD rendering in Octree-GS (Ren et al., 2024), enabling real-time control over detail levels through spatial subdivision.
-3. Preliminaries
-We briefly review 3D Gaussian Splatting and describe its memory usage characteristics. We then introduce the key data structures we used: Gaussian hierarchies and Sequential Point Trees.
-3.1. 3D Gaussian Splatting
-3D Gaussian Splatting (Kerbl et al., 2023) models a radiance field using a set of spatially distributed Gaussians, each with mean $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, RGB base colors $\mathbf{b}_{i}\in\mathbb{R}^{3}$ and covariance matrices $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$.
-The covariance is parameterized via a diagonal scaling matrix $\mathbf{S}_{i}=\text{diag}(s^{1}_{i},s^{2}_{i},s^{3}_{i})$ and an orthonormal rotation matrix $\mathbf{R}_{i}$:
-$\boldsymbol{\Sigma}_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{\top}\mathbf{R}_{i}^{\top}.$
+Compression-based strategies include attribute quantization via codebooks, pruning of low-impact Gaussians, and adapting the degree of spherical harmonics per primitive (ReducingTheMemoryFootprint; Compressed3DGaussianSplatting; LightGaussian; MiniSplatting; HierarchicalCompression; RadSplat; FLoD).
+Scaffold-GS (ScaffoldGS) uses latent vectors anchored to reference Gaussians, with an MLP generating associated Gaussians at render time.
+Octree-GS (OctreeGS) extends this concept to hierarchical LoD rendering, enabling real-time control over detail levels through spatial subdivision.
+3 Preliminaries
+Here, we briefly review 3DGS and how Gaussian hierarchies enable adaptive level-of-detail rendering.
+3.1 Hierarchical 3D Gaussian Splatting
+3DGS (3DGS) models a radiance field using a set of spatially distributed Gaussians, each with mean $\bm{\mu}_{i}\in\mathbb{R}^{3}$, RGB base colors $\mathbf{b}_{i}\in\mathbb{R}^{3}$ and covariance matrices $\bm{\Sigma}_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{\top}\mathbf{R}_{i}^{\top}$, which are parameterized via a diagonal scaling matrix $\mathbf{S}_{i}=\text{diag}(s^{1}_{i},s^{2}_{i},s^{3}_{i})$ and an orthonormal rotation matrix $\mathbf{R}_{i}$.
 Each Gaussian also stores an opacity $\sigma_{i}$ and view-dependent color, modeled using spherical harmonics (SH) coefficients $\mathbf{f}_{i}^{d}$.
@@
 For rendering, all $N$ Gaussians are sorted by distance to the camera and a discrete approximation of the volume rendering equation is evaluated for every pixel $\mathbf{x}$ with corresponding view direction $\mathbf{v}$:
-$\mathbf{C}(\mathbf{x})=\sum_{i=1}^{N}\mathbf{c}_{i}(\mathbf{v})\alpha_{i}(\mathbf{x})\prod_{j=1}^{i-1}(1-\alpha_{j}(\mathbf{x})),$
+$\mathbf{C}(\mathbf{x})=\sum_{i=1}^{N}\mathbf{c}_{i}(\mathbf{v})\alpha_{i}(\mathbf{x})\prod_{j=1}^{i-1}(1-\alpha_{j}(\mathbf{x})),$ (1)
 where $\alpha_{j}$ is the opacity of the $j$-th Gaussian along the view ray:
-$\alpha_{j}(\mathbf{x})=\sigma_{j}e^{-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu^{\prime}}_{j})\boldsymbol{\Sigma^{\prime}}(\mathbf{x}-\boldsymbol{\mu^{\prime}}_{j})^{T}}.$
-Here $\boldsymbol{\mu^{\prime}}$ and $\boldsymbol{\Sigma^{\prime}}$ denote the projected 2D mean and covariance on the image plane, obtained by applying an affine approximation of the projective transform (Zwicker et al., 2001).
-The rendering pipeline is fully differentiable, enabling end-to-end optimization with ADAM using a combination of $\mathcal{L}_{1}$ and SSIM losses.
-3.2. Memory in 3DGS
-Standard 3DGS pipelines store the full set of per-Gaussian attributes, training images, and optimizer state in GPU memory (VRAM).
-Figure 15 summarises the amount of memory that the various components of a single Gaussian require.
+$\alpha_{j}(\mathbf{x})=\sigma_{j}e^{-\frac{1}{2}(\mathbf{x}-\bm{\mu^{\prime}}_{j})\bm{\Sigma^{\prime}}(\mathbf{x}-\bm{\mu^{\prime}}_{j})^{T}}.$ (2)
+Here $\bm{\mu^{\prime}}$ and $\bm{\Sigma^{\prime}}$ denote the projected 2D mean and covariance on the image plane, obtained by applying an affine approximation of the projective transform (zwicker2001EWA).
+3DGS Memory
+Standard 3DGS pipelines store the full set of per-Gaussian attributes, training images, and optimizer state in GPU memory (VRAM); see Figure A.1 for a detailed per-Gaussian breakdown.
 Additional temporary allocations occur during forward and backward passes (e.g., for sorting and gradient accumulation).
 This overhead varies with the scale of the Gaussians and effectiveness of culling strategies, but can be roughly upper-bounded by around 800 bytes per Gaussian in practice.
-This limits training to roughly $500$K Gaussians per GB of VRAM, imposing strict constraints on the detail and extent of reconstructions.
-3.3. Gaussian Hierarchies and Sequential Point Trees
-Gaussian hierarchies, introduced in (Kerbl et al., 2024), recursively merge nearby Gaussians into a tree, where each non-leaf node approximates its children, and leaves correspond to the original Gaussians.
+This limits typical training to roughly 500 000500\,000 Gaussians per GB of GPU memory, imposing strict constraints on the detail and extent of reconstructions.
+Gaussian hierarchies
+as introduced in H-3DGS (Hierarchical), recursively merge nearby Gaussians into a tree, where each non-leaf node approximates its children, and leaves correspond to the original Gaussians.
 A cut is defined by a condition $c_{\text{hier}}(i,\text{cam})$ evaluated in a breadth-first search (BFS).
@@
 A proper cut set includes no parents or children of any included node and thus provides a view-adaptive LoD representation.
-The cut condition used in (Kerbl et al., 2024) is:
-$c_{\text{hier}}(i,\text{cam})=\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i),\quad m_{d}(i)=\frac{T}{\max_{j}s_{i}^{j}},$
+The cut condition used in Hierarchical is a simple cutoff to the camera distance:
+$c_{\text{hier}}(i,\text{cam})=\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i),\quad m_{d}(i)=\frac{T}{\max_{j}s_{i}^{j}},$ (3)
 where $\mathbf{p}_{cam}$ is the camera position, $T$ is a global LoD threshold, and $m_{d}(i)$ is the minimum acceptable distance for viewing Gaussian $i$.
-This ensures that if $i$ is in the cut set, its parent failed the condition:
-$m_{d}(\text{parent}(i))>\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$
-Sequential Point Trees (SPTs) (Dachsbacher et al., 2003), originally developed for point cloud LoD rendering, can be adapted for Gaussians. They enforce a more constrained cut condition:
-$c_{\text{SPT}}(i,\text{cam})=m_{d}(\text{parent}(i))>\left\|\boldsymbol{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$
-This condition is evaluated for all Gaussians in parallel, using the shared root-camera distance $\left\|\boldsymbol{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$.
+The BFS ensures that if $i$ is in the cut set, parent($i$) failed the condition: $m_{d}(\text{parent}(i))>\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$
+4 Method
+To train models that exceed GPU memory limits, all Gaussian attributes are stored in CPU RAM and streamed to the GPU on demand for each training view according to the LoD hierarchy.
+To accelerate the hierarchy cut, we store a copy of only the tree structure in VRAM, where larger subtrees are replaced by Sequential Point Trees (SPTs), forming a hierarchical SPT (HSPT).
+To minimize costly transfers between RAM and VRAM, we track which SPTs currently reside in GPU memory and at which detail.
+Only if an SPT is not present in this GPU cache at a similar level of detail, will it be loaded from RAM.
+Densification is performed on the CPU by adding new leaf nodes to the hierarchy and respawning low-opacity leaf nodes.
+The densified hierarchy is then converted back to an HSPT and transferred to the GPU for a new round of training iterations.
+Appendix A.2 includes more details on initialization and training.
+An overview of our training and densification process can be found in Figure 8 and pseudocodes in Appendix A.8.
+4.1 Sequential Point Trees for Gaussian Splatting
+Figure 1: An SPT and Gaussian hierarchy represent the same 5 Gaussians at varying levels of detail shown with four possible hierarchy cuts in red.
+Vertical lines in the SPT show the binary search result and horizontal lines the distance cut.
+Sequential Point Trees (SPT) were originally developed for point cloud LoD rendering, but can be trivially extended to ellipsoids and Gaussians.
+They enforce a more constrained but more efficient cut condition than Hierarchical:
+$c_{\text{SPT}}(i,\text{cam})=m_{d}(\text{parent}(i))>\left\|\bm{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$ (4)
+This condition is evaluated for all Gaussians in parallel, using the shared root-camera distance $\left\|\bm{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$.
 It requires storing only sorted pairs $\big(m_{d}(i),m_{d}(\text{parent}(i))\big)$, significantly reducing memory compared to full Gaussian hierarchies.
-Note that cuts are guaranteed to be proper, with nodes where $m_{d}(i)<m_{d}(\text{parent}(i))$ never being selected for the cut.
 To optimize cuts, Gaussians are sorted by $m_{d}(\text{parent}(i))$ in descending order.
 A binary search determines the cutoff index $N$, above which Gaussians are too fine to be rendered.
-The downside of the approach is that all Gaussians in the same SPT share the same level of detail dictated by the distance from the camera to the root node.
-This can lead to Gaussians with $m_{d}(i)>\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}$ being rendered, even though they would be too coarse for the current view.
-To counteract this issue, we define:
-$M_{d}(i)=m_{d}(i)+\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2},$
-as a conservative minimum distance function.
-By the triangle inequality, selecting Gaussians satisfying $M_{d}(i)\leq\left\|\boldsymbol{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$ guarantees $m_{d}(i)\leq\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}$.
+Note that cuts are guaranteed to be proper, with nodes where $m_{d}(\text{parent}(i))>m_{d}(i)$ never being selected for the cut.
+Evidently, the level of detail of all Gaussians in the SPT is dictated by the camera’s distance to its root node.
+This can lead to Gaussians with $m_{d}(i)>\left\|\bm{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}$ being rendered, even though they would be too coarse for the current view.
+To counteract this issue, we define: $M_{d}(i)=m_{d}(i)+\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2},$ as a conservative minimum distance function.
+By the triangle inequality, selecting Gaussians satisfying $M_{d}(i)\leq\left\|\bm{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$ guarantees $m_{d}(i)\leq\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}$.
 In turn, this means that Gaussians that are further away from the camera than the root node will be selected at a higher level of detail.
@@
 Figure 1 visualizes both hierarchy types and their LoD cuts for a toy example.
-Figure 2.
-A Sequential Point Tree and Gaussian hierarchy represent the same 5 Gaussians at varying levels of detail shown with four possible hierarchy cuts in red.
-Horizontal lines in the SPT show the binary search result and horizontal lines the distance cut.
-[…]
-[SPT vs Gaussian hierarchy]Sequential Point Tree and Gaussian hierarchy datastructure for a set of 9 Gaussians along with various levels of detail and their cuts
-4. Method
-An overview of the training and densification process can be found in Figure 7, where we enumerate the individual steps of our training and densification process: To train models exceeding GPU memory limits, all Gaussian attributes are stored in CPU RAM and streamed to the GPU on demand for each training view.
-To accelerate the hierarchy cut, we store a copy of only the tree structure in VRAM, where larger subtrees have been replaced by SPTs, forming a hierarchical SPT (HSPT).
-To minimize costly transfers between RAM and VRAM, we keep track of which SPTs are currently loaded and at which detail, and cache them in GPU memory.
-Only if an SPT is not present in this GPU cache at a similar level of detail, will it be loaded from RAM.
-Densification occurs on the CPU by adding new leaf nodes to the hierarchy and respawning low-opacity leaf nodes.
-The densified hierarchy is then converted back to an HSPT and transferred to the GPU for a new round of training iterations.
-4.1. Initialization
-Following Kerbl et al.
-(2024), we initialize the Gaussian model from a sparse point cloud, augmented with skybox points.
-This initial representation is small enough to fully reside in GPU memory and is trained for 100 000100000100\,000100 000 iterations with densification disabled.
-The goal of this phase is to establish a stable global scene structure before hierarchy construction.
-After this initial optimization, we build a binary Gaussian hierarchy, with the trained Gaussians as leaf nodes and parent nodes representing merged approximations of their children.
-4.2. Gradient Propagation
-To update the properties of Gaussians, including base color, SH coefficients, position, and covariance, we follow the standard 3DGS error propagation (Kerbl et al., 2023).
-However, we operate only on the cut set chosen for each training view.
-Thus, gradients may, in general, be propagated to Gaussians in the middle of the hierarchy.
-4.3. Densification
-Densifying a Level-of-Detail representation presents a unique challenge: the underlying hierarchical structure must evolve continuously during training.
+4.2 Densification
+Figure 2: Example of densifying and respawning leaf nodes.
+Densifying an LoD representation presents a unique challenge, as the hierarchical structure must evolve continuously during training.
 Prior works circumvent this issue by constructing LoD hierarchies only after chunk-level training and densification are complete.
-We take inspiration from 3DGS-MCMC (Kheradmand et al., 2024), which selects Gaussians probabilistically (weighted by opacity) to be ‘split’, replacing it by two new Gaussians which together should appear similarly to the original split Gaussian.
-Noteably, this resembles the way a parent node in a Gaussian hierarchy approximates its children.
-Therefore, we adopt this approach and ‘spawn’ two new child nodes for a leaf instead of splitting a Gaussian, increasing the size of the hierarchy with minimal artifacts.
-Instead of pruning, Kheradmand et al.
-(2024) declares Gaussians below a certain opacity threshold as ‘dead’, and respawns them at the position of a high-opacity Gaussian.
+We take inspiration from 3DGS-MCMC (3DGSasMCMC), which ‘splits’ Gaussians, replacing them with two new Gaussians which together should appear similarly to the original.
+Notably, this mirrors how a parent node in a Gaussian hierarchy approximates its children.
+Therefore, we adopt this approach and instead ‘spawn’ two new child nodes for a Gaussian, increasing the size of the hierarchy with minimal artifacts.
+Instead of pruning, 3DGSasMCMC declares Gaussians below a certain opacity threshold as ‘dead’, and respawns them at the position of a high-opacity Gaussian.
 We propose a similar strategy: when a leaf node dies, its parent is replaced by its sibling node; the dead leaf node and its parent are then respawned as children to another node, which is selected to be densified.
@@
 Together, they ensure that the hierarchy can be expanded during training in a stable manner and rebalanced as required.
-Avoiding destructive resets, as advocated by Kerbl et al.
-(2023), is particularly important in our setting, as many Gaussians will no longer conform to the hierarchical structure after opacity is restored to normal.
-Figure 3. Example of densifying and respawning leaf nodes.
-4.4. The hierarchical SPT datastructure
+While 3DGSasMCMC choose Gaussians to densify using a random selection weighted by opacity, we employ the strategy from Hierarchical, which selects Gaussians for densification based on their maximal screen-space gradient.
+Avoiding destructive opacity resets, as advocated by 3DGS, is particularly important in our setting, as many Gaussians will no longer conform to the hierarchical structure after opacity is restored to normal.
+Further details on our densification method can be found in Appendix A.7.
+4.3 The hierarchical SPT datastructure
+This section covers previous approaches to LoD selection and why a new datastructure—the hierarchical SPT—is necessary for robust and efficient training.
 BFS
@@
 Parallel Cut
-To enable GPU-accelerated cuts, Kerbl et al. (2024) evaluate the cut condition in parallel for each Gaussian:
-$\Big{(}m_{d}(i)<\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}\Big{)}\land\Big{(}m_{d}(parent(i))\geq\left\|\boldsymbol{\mu}_{\text{parent}(i)}-\mathbf{p}_{cam}\right\|_{2}\Big{)},$
+To enable GPU-accelerated cuts, Hierarchical evaluate the cut condition in parallel for each Gaussian:
+$\Big(m_{d}(i)<\left\|\bm{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}\Big)\land\Big(m_{d}(parent(i))\geq\left\|\bm{\mu}_{\text{parent}(i)}-\mathbf{p}_{cam}\right\|_{2}\Big),$ (5)
 where any Gaussian that is sufficiently small at its current camera distance and whose parent is too large to be rendered, should be part of the cut set.
-This produces a proper cut under the assumption that child Gaussians always have a smaller minimal distance than their parents:
-$\forall i:m_{d}(i)<m_{d}(parent(i)).$
-This is generally valid when hierarchies are constructed after training, since parent Gaussians represent coarser approximations of their children.
-However, when the hierarchy is modified during training and densification, optimization can cause this assumption to break—leading to invalid cut sets and degenerate hierarchies that worsen over time.
+This produces a proper cut under the assumption that child Gaussians always have a smaller minimal distance than their parents (i.e.
+the heap condition is fulfilled): $\forall i:m_{d}(i)<m_{d}(parent(i)).$ This is generally valid when hierarchies are constructed after training, since parent Gaussians represent coarser approximations of their children.
+However, when the hierarchy is modified during training and densification, optimization can break the heap condition—leading to invalid cut sets and degenerate hierarchies that worsen over time.
 Hierarchical SPT (HSPT)
-We present the hierarchical SPT data structure, which combines the benefits of both approaches. To construct it, we cut it using a BFS on the condition
-$c_{HSPT}(i)=s_{i}^{1}\cdot s_{i}^{2}\cdot s_{i}^{3}<\texttt{size}$
-with volume threshold size.
+Figure 3: A Gaussian hierarchy is converted to an HSPT by cutting according to Gaussian volume and converting sufficiently large subtrees to SPTs. The HSPT can then be cut in a 2-step process.
+We present the hierarchical SPT data structure, which combines the benefits of both approaches.
+To construct it from a Gaussian hierarchy, we cut it using a BFS on the condition $c_{\text{HSPT}}(i)=s_{i}^{1}\cdot s_{i}^{2}\cdot s_{i}^{3}<\texttt{size}$ with volume threshold size.
 The resulting cut set $\mathbb{C}_{\text{HSPT}}$ partitions the hierarchy into the upper hierarchy, which includes all Gaussians with volume greater than size, and the lower hierarchy, consisting of the subtrees rooted at the nodes in the cut set.
@@
 The construction and cutting process of an HSPT is illustrated in Figure 3.
-The SPTs used to render a MatrixCity-Scale frame are shown in Figure 11, and Figure 10 demonstrates the smooth LoD transitions enabled by the hierarchical structure.
-Figure 4. A Gaussian hierarchy is converted to an HSPT by cutting according to Gaussian volume and converting sufficiently large subtrees to SPTs. The HSPT can then be cut in a 2-step process.
-Rebuilding the hierarchical SPT for every training iteration would eliminate any performance benefit.
+Figure 10 and Figure 15 show example frames with respectively SPTs highlighted and different levels of detail.
+Rebuilding the HSPT for every training iteration would eliminate any performance benefit.
 Instead, we exploit the fact that the minimum distance $m_{d}$ evolves slowly during optimization and thus only needs to be updated infrequently.
@@
 This infrequent recomputation allows us to use a more accurate—albeit more expensive—minimum distance metric than the inverse of maximal scale. Specifically, we define:
-$m_{d}^{\prime}(i)=\frac{T}{\sqrt{s_{i}^{1}\cdot s_{i}^{2}+s_{i}^{1}\cdot s_{i}^{3}+s_{i}^{2}\cdot s_{i}^{3}}},$
+$m_{d}^{\prime}(i)=\frac{T}{\sqrt{s_{i}^{1}\cdot s_{i}^{2}+s_{i}^{1}\cdot s_{i}^{3}+s_{i}^{2}\cdot s_{i}^{3}}},$ (6)
 which corresponds to the inverse square root of the surface area of the Gaussian ellipsoid (up to a constant factor).
@@
 To this end, we frustum cull every node considered in the BFS by checking if a sphere around the Gaussians with radius equal to $(3\cdot\max_{j}s_{i}^{j})$ intersects with the view frustum.
-Using the Gaussian scale as a proxy for the extent of its entire subtree is not perfectly accurate, but comparisons to using a full bounding sphere hierarchy showed no discernable difference in practice.
+Using the Gaussian scale as a proxy for the extent of its entire subtree is not perfectly accurate, but comparisons to using a full bounding sphere hierarchy showed no discernable difference in our experiments.
 It should be noted that Gaussians are implicitly frustum culled during rasterization, but this early culling accelerates the cutting procedure and significantly decreases the number of Gaussians that need to be loaded from RAM, as can be seen in Figure 9.
-4.5. Caching on the GPU
+4.4 Caching on the GPU
+Figure 4: Gaussians required for the current training view are assembled from three sources: the upper tree, newly loaded SPT cuts from RAM, and cache hits.
+After optimization, newly accessed SPTs are added to the GPU cache.
 Loading Gaussian data from RAM is a costly operation that can become a significant bottleneck during large-scale training.
 To mitigate this, we maintain a GPU-resident cache of Gaussians that are likely to be reused across consecutive training views.
-However, checking the cache for every individual Gaussian would introduce nontrivial overhead.
+However, checking the cache for every individual Gaussian would introduce non-trivial overhead.
 Once again, SPTs offer an efficient alternative.
-Rather than caching individual Gaussians, we store the Gaussians from SPT cuts along with the cached distance from the camera to the root of each SPT, denoted $\bar{d}^{j}$ for the $j$th SPT.
-During rendering, when the upper hierarchy is cut and the required SPTs identified, we compute $d^{j}=\|\boldsymbol{\mu}_{\text{root}(j)}-\mathbf{p}_{\text{cam}}\|_{2}$ and check whether a matching cut is already cached.
-This check is performed using a simple distance ratio tolerance:
-$D_{\text{min}}\leq\frac{d^{j}}{\bar{d}^{j}}\leq D_{\text{max}}.$
+Rather than caching individual Gaussians, we store the Gaussians from SPT cuts along with the cached distance from the camera to the root of each SPT, denoted $\bar{d}^{j}$ for the $j$-th SPT.
+During rendering, when the upper hierarchy is cut and the required SPTs identified, we compute $d^{j}=\|\bm{\mu}_{\text{root}(j)}-\mathbf{p}_{\text{cam}}\|_{2}$ and check whether a matching cut is cached, using a simple distance ratio tolerance:
+$D_{\text{min}}\leq\frac{d^{j}}{\bar{d}^{j}}\leq D_{\text{max}}.$ (7)
 Here, $D_{\text{max}}$ defines the allowable range for using coarser-than-optimal detail, while $D_{\text{min}}$ limits how much finer detail can be tolerated.
@@
 After each training iteration, newly loaded SPTs are added to the cache.
-To bound VRAM usage, the we use a least-recently-used (LRU) write-back policy.
+To bound VRAM usage, we use a least-recently-used (LRU) write-back policy.
 When a memory threshold is exceeded, entries are written back to RAM.
-Additionally, to prevent overfitting to persistent cache entries, the entire cache is flushed every 1 00010001\,0001 000 iterations.
+Additionally, to prevent overfitting to persistent cache entries, the entire cache is flushed every 1 0001\,000 iterations.
 Figure 4 illustrates the caching process across two frames.
-Figure 5.
-Caching strategy overview.
-Gaussians required for the current training view are assembled from three sources: the upper tree, newly loaded SPT cuts from RAM, and cache hits.
-After optimization, newly accessed SPTs are added to the GPU cache.
 View Selection
 In large-scale scenes, the GPU cache typically covers only a small fraction of the overall geometry, leading to sparse cache hits.
-To improve cache utilization, we prioritize spatial locality by selecting successive training views that are geographically close to the current one, thereby maximizing Gaussian reuse.
-To this end, we precompute a $k$-nearest-neighbor graph over all training view positions, where edge weights $w_{ij}$ correspond to the Euclidean distance between views $i$ and $j$.
-The next training view $j$ is then sampled from the $k$-nearest neighbors of the current view $i$ according to the distribution:
-$\mathbb{P}(j\mid i)\propto\frac{1}{w_{ij}+W},$
-where $W$ is a constant, controlling the degree of exploration.
-However, care must be taken when modifying the order of training views, as deviating from uniform random sampling may introduce bias.
-Thus, we inject a random view every 12 iterations, which we find sufficient to preserve generalization performance.
-4.6. Memory Layout
-Figure 6 illustrates the peak memory usage for a single training iteration of a 60-million-Gaussian hierarchy on the MatrixCity-Scale dataset.
+To improve cache utilization, we prioritize spatial locality by selecting successive training views close to the current one, maximizing Gaussian reuse.
+To this end, we precompute a $k$-nearest-neighbour graph over all training view positions, where edge weights $w_{ij}$ correspond to the Euclidean distance between views $i$ and $j$.
+The next training view $j$ is then sampled from the $k$-nearest neighbours of the current view $i$ according to the distribution: $\mathbb{P}(j\mid i)\propto\frac{1}{w_{ij}+W},$ where $W$ is a normalization constant that also controls the degree of exploration.
+However, care must be taken when modifying the order of training views, as deviating from uniform random sampling may introduce training bias.
+To counteract this, we inject a randomly selected view every 128 iterations, which we find sufficient to preserve generalization performance.
+4.5 Memory Layout
+Figure 5: Peak memory consumption of CPU and GPU for a training iteration on MC-smaller-city+ with 60 million Gaussians.
+Figure 5 illustrates the peak memory usage for a single training iteration of a 60-million-Gaussian hierarchy on the MC-smaller-city+ dataset.
 The output frame from this iteration is shown in Figure 13.
@@
 Therefore, minimizing the number of active Gaussians is critical for staying within the VRAM budget.
-The remaining GPU memory usage consists of auxiliary data, including cache management, hierarchy cut tracking, training and ground-truth images, as well as PyTorch overhead.
-Figure 15 provides a breakdown of memory usage for Gaussian properties and hierarchy data, and how these are distributed across CPU and GPU.
-Figure 6. Peak memory consumption of CPU and GPU for a training iteration on BigCity-Scale with 60 million Gaussians.
-5. Evaluation
-Choice of dataset
-Our method is designed to enable seamless training and rendering on large-scale scenes comprising thousands of views captured at vastly different scales.
-Unfortunately, most existing datasets of sufficient size contain either street-level or aerial views—but not both.
-Additionally, at this scale and density, current Structure-from-Motion (SfM) pipelines often fail to reliably recover camera poses and sparse point clouds.
-To overcome these challenges, we use the synthetic MatrixCity dataset (Li et al., 2023), which provides dense multi-scale views along with ground-truth camera poses and sparse geometry.
-For our main benchmark, MatrixCity-Scale (MC-Scale), we aggregate 11 466 street-view images, 3 208 aerial views, and 533 additional high-altitude views we generate manually.
-We evaluate reconstruction quality on a separate set of 1 694 test images.
-The diversity in image scale is visualized in Figure 8.
-To validate our method on real-world data, we also include results on the largest public dataset from Kerbl et al.
-(2024), the Campus scene.
-Choice of Comparison
-Selecting appropriate baselines is challenging.
-Standard 3DGS variants perform poorly at large scales, while most divide-and-conquer approaches—such as Hierarchical 3DGS, CityGS, and HorizonGS—assume access to 80 GB GPUs.
-Despite access to multiple clusters, we were limited to 40 GB A100s, highlighting the hardware demands of these methods.
-To our knowledge, only Hierarchical 3DGS has trained successfully on the Campus dataset, and no prior work has demonstrated large-scale training on MatrixCity with combined aerial and street views.
-Due to resource constraints, we were unable to train CityGS on either dataset and adopt Hierarchical 3DGS as our main baseline.
-For additional reference, we evaluate 3DGS-MCMC, modified to support large-scale training by limiting densification, streaming images from disk, and reducing the SH degree.
-Most chunk-based methods, including Kerbl et al.
-(2024), require feature correspondences, which are not provided in MatrixCity.
-To enable comparison, we reconstructed a sparse point cloud using COLMAP and the provided camera poses.
-This less stable point cloud (visualized in Figure 8) was used for initialization.
-On the Campus scene, we report results for the official test split (Hierarchical single) and the full pretrained model.
-Consequently, we evaluate reconstruction quality on every 100th training view.
-We train a smaller model (38M) and a larger (80M) model.
-Rendering Hierarchical 3DGS required reducing SH degree to 1—even on 40 GB VRAM; ours also uses SH degree 1.
-Given the limitations of numerical metrics in these settings, we strongly encourage visual comparison via the supplementary videos, which in our opinion favour Hierarchical 3DGS.
-Except for the results of Hierarchical 3DGS, all training and rendering was done using an RTX 3090 GPU.
-5.1. Results
-Quantitative comparisons are shown in Table 1.
-A LoD of Gaussians successfully reconstructs the full MC-Scale scene, where other methods struggle due to the scene’s size and diverse camera viewpoints.
-Notably, Kerbl et al.
-(2024) reconstructs individual chunks accurately, but the merging process introduces ”floaters”—isolated Gaussians not associated with meaningful geometry—which obscure most test views (see Figure 13 and 14).
-Similarly, 3DGS-MCMC suffers from aggressive opacity and scale-based pruning, causing a large fraction of Gaussians to vanish entirely.
-On the Campus dataset, Hierarchical 3DGS performs well, as its chunk-based approach aligns with the dataset’s single, street-level trajectory.
-Their single-chunk evaluation shows strong metrics.
-However, when evaluating their full model with reduced SH degree (to fit within 40 GB VRAM), we observe a clear drop in numerical quality.
-That said, the visual results remain competitive—see the supplemental video.
-In our case, both low- and high-resolution variants achieve similar quantitative scores despite visible differences, due to Hierarchical 3DGS evaluating on training views.
-For context, Hierarchical 3DGS splits the MC-Scale scene into 94 chunks, each trained for 45 000 iterations, totaling over 4.23 million steps and nearly four full days of compute time on our machine.
-By contrast, our approach trains the full scene end-to-end in just 250 000 iterations, completing in approximately 15 hours.
-Memory
-A major limitation of chunk-based methods is the memory overhead during rendering.
-Although it also implements a form of GPU streaming, we were unable to run the viewer provided by Kerbl et al.
-(2024) on either dataset using the 40 GB VRAM A100, which is consistent with the limitations they report.
-In contrast, our method renders interactive flythroughs of MatrixCityScale and Campus (38M), see supplemental, while only requiring 8 GB of GPU memory (and 16 GB for Campus (80M)).
-Table 1.
-Full Novel View Synthesis results.
-* use of the suboptimal COLMAP version for MC-Scale; adjusted SH degree to match ours and reduce memory to enable rendering, use of higher SH degree and only single chunk, reproduced from (Kerbl et al., 2024).
-MC-Scale (15k images) Campus (22k images)
+The remaining GPU memory usage consists of auxiliary data, including cache management, hierarchy cut tracking, training and ground-truth images, as well as general PyTorch overhead.
+Figure A.1 provides a breakdown of all data associated with a single Gaussian and how they are distributed across CPU and GPU.
+5 Evaluation
+Our method is designed to enable seamless training and rendering on ultra-large-scale scenes comprising tens of thousands of views captured at vastly different scales.
+Unfortunately, most datasets of sufficient size contain either street-level or aerial views—but not both.
+To overcome these challenges, we turn to the synthetic MatrixCity dataset (MatrixCityDataset), which provides dense multi-scale views along with ground-truth camera poses and sparse geometry.
+We also include results on the real-world street-level datasets from H-3DGS.
+As baseline, we compare to recent large-scale 3DGS methods CityGaussian (CityGaussian), H-3DGS (Hierarchical), HorizonGS (HorizonGS) and OctreeGS (OctreeGS).
+We note that HorizonGS is not evaluated on the H-3DGS dataset, since it requires aerial views for training.
+Further results and additional details on the datasets and baselines and are included in A.4.
+Evaluations were run on a single H200 GPU with 141 GB VRAM, while all of our models were trained on a single RTX 3090 graphics card with 24 GB VRAM.
+The contribution of our components is ablated in Appendix A.3.
+5.1 Results
+Qualitative comparisons are shown in Figure 6 and quantitative comparisons in Tables 1 and 2.
+The outputs of H-3DGS and HorizonGS show significant floating and ghosting artifacts.
+While H-3DGS performs well on individual chunks of MC-smaller-city+, oversized floaters survive the merging procedure and obscure the majority of test images (cf.
+Figure 11, 13 and 17).
+The merging procedure of CityGaussian (CityGaussian), which is designed for aerial-only datasets, discards most of the trained Gaussians to avoid chunk artifacts, leading to mostly artifact-free views, but significantly blurry results.
+A LoD of Gaussians reconstructs novel views with a high degree of detail while avoiding chunk-based artifacts (cf.
+Figure 16), resulting in a considerable lead in quality metrics.
+Seamless training enables faster convergence during training, reducing the required number of training iterations considerably over divide-and-conquer based methods.
+Table 1: MatrixCity novel view synthesis results.
+Results with * use the MC-smaller-city+ dataset due to memory constraints.
+Results with require suboptimal COLMAP initialization instead of the dataset’s provided point cloud.
+VRAM usage is measured while rendering the test images.
+Methods are considered out-of-memory (OOM) if they exceed 141GB of VRAM during rendering or training.
+MC-small-city+ (42.2k images) / *MC-smaller-city+ (15.1k images)
 […]
-Ours* 20.63 0.668 0.282 40M 22.83 0.713 0.248 38M
-Hierarchical*,‡ 13.77 0.519 0.559 82M 17.76 0.601 0.424 80M
-Ours 21.73 0.712 0.213 60M 22.86 0.725 0.237 80M
-Hierarchical single 24.61 0.807 0.331 -
-3DGS-MCMC 13.16 0.4481 0.586 6M 15.11 0.600 0.580 6M
-Table 2.
-Ablations.
-Average frame times for rendering camera paths and average iteration times during training with and without caching Gaussians.
-Then, average times to cut the hierarchy during rendering with the normal BFS approach and HSPT.
-80M Campus was trained longer than 38M.
-MC-Scale Campus
-38M 80M
-Render Full 48.1 ms 47.1 ms 83.2 ms
-Render w/o Cache 119.4 ms 92.6 ms 222.3ms
-Render w/o Frustum Culling 52.3 ms 38.3 ms 110.2 ms
-Train Full 156 ms 205 ms
-Train w/o Cache 471 ms 244 ms
-Train w/o Frustum Culling 685 ms 312 ms
-HSPT Cut (Render) 31.9 ms 31.3 ms 36.5 ms
-BFS Cut (Render) 47.8 ms 40.0 ms 53.7 ms
-5.2. Ablations
-We assess the contribution of key components through an ablation study, using recorded camera paths across all scenes (see supplemental video).
+CityGaussian 19.78 0.650 0.475 2.4 GB 1.11M 2.7M
+HorizonGS 12.06 0.521 0.544 20.6 GB 2.5M 50.3M
+H-3DGS OOM OOM OOM >>141GB OOM OOM
+OctreeGS OOM OOM OOM >>141GB OOM OOM
+Ours 21.59 0.711 0.396 25.0GB 600k 136.7M
+HorizonGS* 15.28 0.601 0.435 13.91 1.2M 34.1M
+OctreeGS* OOM OOM OOM >>141GB OOM OOM
+Ours* 21.71 0.710 0.370 17.7GB 310k 60M
+H-3DGS* 14.57 0.516 0.563 97GB 4M 83.1M
+Ours* 20.59 0.658 0.468 21.7GB 310k 24.7M
+Table 2: Hierarchical 3DGS dataset novel view synthesis results.
+Small City (5.8k images) Campus (22.0k images)
+[…]
+OctreeGS 19.20 0.601 0.498 OOM OOM OOM
+CityGS 21.40 0.692 0.383 19.47 0.645 0.531
+H-3DGS (leaves) 24.35 0.788 0.273 17.84 0.6101 0.4667
+Ours 24.61 0.774 0.297 21.85 0.707 0.456
+(a) Ground Truth
+(b) Ours
+(c) CityGaussian
+(d) Hierarch. 3DGS
+(e) HorizonGS
+Figure 6: Qualitative comparison of our method and SOTA methods (CityGaussian; Hierarchical; OctreeGS) on the MC-small-city+ and campus datasets.
+Rendering
+Our level-of-detail and caching strategy can also be applied to efficiently render the finished models: As supplemental material, we include fly-through videos of all scenes rendered on an RTX 3090 GPU at this anonymized link: https://www.youtube.com/playlist?list=PLocO6QqSKZ4YN7DyuC5gWEeVciSV6RdAk.
+Table 1 demonstrates effective VRAM reduction during rendering of test images in comparison with other large-scale 3DGS methods.
+Comparisons of rendering VRAM with 3DGS (3DGS) and gsplat (Gsplat) are included in Appendix A.4.1.
+6 Discussion and Outlook
+A LoD of Gaussians enables seamless training and rendering of ultra-large 3DGS models on consumer hardware.
+By storing Gaussian data in external memory and streaming it on demand, our method avoids the pitfalls of chunk-based pipelines.
+The hierarchical SPT structure accelerates LoD selection and remains robust to ongoing training changes.
+Combined with caching and view selection, our system significantly reduces out-of-core overhead.
+Together, these components enable efficient reconstruction and rendering at scale, demonstrated on the MC-small-city+ dataset.
+Limitations and Future Work
+A primary limitation of our method lies in initialization.
+Accurate estimation of camera poses and sparse point clouds remains challenging at large scale, particularly for real-world datasets with sparse or inconsistent coverage.
+Future work could explore improved initialization strategies or joint optimization of poses and Gaussians during early training.
+Our system also requires roughly 1 GB of RAM per million Gaussians, which—while more efficient than prior methods—still constrains scalability.
+Loading from disk is feasible in our experiments, though at the cost of a $10\times$ slowdown.
+Further, while our method in general greatly reduces the number of training iterations necessary, individual training iterations can take a lot longer than standard 3DGS due to the loading overhead.
+The level-of-detail system makes our method robust against vast differences in view distance, but when these differences are not present, like in single-height aerial datasets as evaluated in Appendix A.4, it becomes a burden instead, making more straight-forward divide-and-conquer training the better choice.
+Interactive rendering could be further optimized by avoiding per-frame hierarchy cut recomputation and enabling asynchronous streaming.
+While frustum culling reduces memory load in most views, it is ineffective when the entire scene falls inside the frustum.
+Occlusion culling could address this by skipping entire SPTs before loading.
+Overall, we believe that out-of-core 3D Gaussian Splatting is a promising direction for scaling radiance field methods to city-scale and beyond, without the need for specialized hardware.
+Reproducibility Statement
+To ensure reproducibility of our results, we have appended our entire codebase along with the config files of our method and baselines required to train each of the scenes.
+Detailed breakdowns of datasets and hyperparameters can be found in Appendix A.4 and A.5.
+Appendix A Appendix
+A.1 Figures
+(a) Divide and Conquer
+(b) Ours
+Figure 7: Comparison between the divide and conquer training process as used in Hierarchical; VastGaussian; CityGaussian; HorizonGS to our training process.
+Colored regions are present in VRAM, training views are drawn in red.
+[…]
+Figure 9: Frustum Culling and LoD selection (left) greatly reduces the number of Gaussians required to render a view (right).
+Figure 10: Hierarchical SPTs enable smooth transitions between detailed (left) and coarse (right) representation.
+(a) Chunk Bleeding in H-3DGS (Hierarchical).
+(b) Chunk Ghosting in CityGaussian (CityGaussian).
+Figure 11: Artifacts caused by the divide-and-conquer strategy on the MC-small-city+ dataset.
+Figure 12: Training image rendered during the iteration depicted in Figure 5.
+Figure 13: Images from H-3DGS on the MatrixCity-Scale scene are obscured by large floaters.
+Figure 14: All Gaussian properties except gradients are always present in CPU RAM, whereas only the currently loaded Gaussians and slim SPT information needs to be stored on GPU.
+Figure 15: SPTs for a frame of MatrixCity rendered in different colors.
+Figure 16: A LoD of Gaussians is able to reconstruct and render huge environments with 150M+ Gaussians with large variations of scale on consumer hardware.
+(a) Image generated during chunk training by H-3DGS.
+(b) The same image generated during training by A LoD of Gaussians.
+Figure 17: Chunk-based optimizations causes floaters to expand outside of the chunk boundary.
+A.2 Initialization and Training
+Following Hierarchical, we initialize the Gaussian model from a sparse point cloud, augmented with skybox points.
+This initial representation is small enough to fully reside in GPU memory and is trained for 100100k iterations without densification.
+The goal of this phase is to establish a stable global scene structure before constructing the hierarchy.
+After this initial optimization, we build a binary Gaussian hierarchy, where the trained Gaussians act as leaf nodes and the parent nodes represent merged approximations of their two children.
+The Gaussian properties–including base color, SH coefficients, position, and covariance–are optimized using the standard loss propagation introduced by 3DGS.
+To fulfill our tight memory requirements, we update only the parameters of the cut set chosen for each training view.
+Thus, gradients may, in general, be propagated to Gaussians in the middle of the hierarchy.
+Assuming good coverage and diversity of training views, these updates will diffuse over the entire hierarchy over the course of training, leading to smooth transitions from the highest to the lowest LoD (see Figure 10).
+A.3 Ablations
+We assess the contribution of key components through an ablation study, using recorded camera paths across all scenes (see supplemental videos).
 Table 3 reports average frame times over these paths.
@@
 For training, we measure average iteration durations over 1 000 steps.
-Here, frustum culling again proves essential, substantially reducing the number of Gaussians loaded and processed per view.
+Here, frustum culling and caching prove essential, substantially reducing the number of Gaussians loaded and processed per view.
+Table 3: Ablations.
+Average frame times for rendering camera paths and average iteration times during training with and without caching Gaussians.
+The final results show the average timings of the hierarchy cut during rendering using our HSPT and the baseline BFS approach.
+For campus, we evaluate two differently sized models with 38M and 80M Gaussians respectively.
+MC-smaller-city+ MC-small-city+ Campus
+60M 136M 38M 80M
+Render Full 48.1 ms 68.3ms 47.1 ms 83.2 ms
+Render w/o Cache 119.4 ms 127.4 ms 92.6 ms 222.3ms
+Render w/o Frustum Culling 52.3 ms 89.3 38.3 ms 110.2 ms
+Train Full 156 ms 174 ms 205 ms
+Train w/o Cache 471 ms 257 ms 244 ms
+Train w/o Frustum Culling 685 ms 626 ms 312 ms
+HSPT Cut (Render) 31.9 ms 45.7ms 31.3 ms 36.5 ms
+BFS Cut (Render) 47.8 ms OOM 40.0 ms 53.7 ms
 View Selection
-On the MatrixCity-Scale dataset (250K iterations, up to 60M Gaussians), enabling view selection reduces the number of Gaussians loaded from RAM per frame from 355K to 219K—a $35\%$ reduction in memory load.
+When training MC-smaller-city+ dataset (250k iterations, up to 60M Gaussians), enabling our guided view selection reduces the number of Gaussians loaded from RAM per frame from 355 770 to 219 043—a $35\%$ reduction in memory transfers.
 Reconstruction quality is unaffected, with PSNR and SSIM improving slightly by 0.04 and 0.05, respectively.
-6. Discussion and Outlook
-A LoD of Gaussians enables seamless training and rendering of ultra-large 3DGS models on consumer hardware.
-By storing Gaussian data in external memory and streaming it on demand, our method avoids the pitfalls of chunk-based pipelines.
-The hierarchical SPT structure accelerates LoD selection and remains robust to ongoing training changes.
-Combined with caching and view selection, our system significantly reduces out-of-core overhead.
-Together, these components enable efficient reconstruction and rendering at scale, demonstrated on the MatrixCity-Scale dataset.
-Limitations and Future Work
-A primary limitation of our method lies in initialization.
-Accurate estimation of camera poses and sparse point clouds remains challenging at large scale, particularly for real-world datasets with sparse or inconsistent coverage.
-Future work could explore improved initialization strategies or joint optimization of poses and Gaussians during early training.
-Our system also requires roughly 1 GB of RAM per million Gaussians, which—while more efficient than prior methods—still constrains scalability.
-Loading from disk is feasible in our experiments, though at the cost of a 10×10$\times$10 × slowdown.
-Preliminary exploration has also shown that given access to sufficient RAM and training time, our method could scale far beyond what is presented here.
-A 150 million Gaussian model has been trained successfully for MatrixCity-Scale on 24 GB of VRAM, though the improvement to image quality was minor.
-We suspect that even more training views are required to properly leverage such a vast number of Gaussians.
-Rendering could be further optimized by avoiding per-frame hierarchy cut recomputation and enabling asynchronous streaming.
-While frustum culling reduces memory load in most views, it is ineffective when the entire scene falls inside the frustum.
-Occlusion culling could address this by skipping entire SPTs before loading.
-Overall, we believe that out-of-core 3D Gaussian Splatting is a promising direction for scaling radiance field methods to city-scale and beyond, without the need for specialized hardware.
+Effectiveness of LOD
+To evaluate the effectiveness of our level-of-detail system, we choose 5 random images from each dataset and compare the resulting quality metrics and number of Gaussians rendered at 50 different levels of details.
+The results are visualized in Figure 18.
+In the vast majority of cases, increasing the LOD level consistently improves rendering quality.
+Overall, we only notice a very small dip in quality when reducing the Gaussian count by 50%, verifying the effectiveness of our continuous LoD structure.
+(a) MC-small-city+ aerial
+(b) MC-small-city+ street
+(c) campus
+Figure 18: Quantitative evaluation of our LoD system for 5 randomly chosen views and 50 different levels of detail.
+A.4 Dataset Details and Further Results
+We include more details on the datasets used in the evaluation and provide additional results for smaller-scale scenes.
+MC-small-city+
+For our main benchmark, MC-small-city+, we aggregate 33 006 street-view images and 7 672 aerial views from the small-city scene of the MatrixCity dataset (MatrixCityDataset) and add 533 additional high-altitude views we generated manually.
+We evaluate reconstruction quality on a separate set of 4 228 test views, according to the test split provided by the dataset.
+The scene is extremely challenging due to its enormous scale, sparse views and wide variation in scale: As such, we use a subset of the entire dataset (15.1k images, covering about a third of the area), denoted as MC-smaller-city+ for baselines that are not able to reconstruct the full dataset.
+For MC-small-city+, we use the camera poses provided by the MatrixCity dataset and convert them to the COLMAP format for Gaussian splatting.
+We merge the provided street and aerial sparse point clouds and randomly downsample them by a factor of $5$ to get more realistic initialization conditions.
+We do not make use of the ground truth depth images provided by the dataset, as we consider this to be an unrealistic advantage.
+Because the dataset does not provide point correspondences required for scaling monocular depth maps, we deactivate depth supervision in our method and baselines for this dataset.
+Figure 19 shows the scene along with the camera distribution.
+While some methods have successfully reconstructed only aerial (CityGaussian; CityGaussianV2; OctreeGS) or only street-level views (CityGaussianV2; NerfXL), training a model that holds up to scrutiny from both perspectives presents a particular challenge.
+Training on close and far views simultaneously–without a proper LoD system like ours in place–significantly degrades visual quality for both, as noted by HorizonGS; CrossView.
+Moreover, such a scenario complicates partitioning the scene into independent chunks.
+Figure 19: The MC-small-city+ dataset spans multiple city blocks, which are supervised by tens of thousands of views, varying widely in scale.
+Hierarchical 3DGS dataset
+H-3DGS Hierarchical did not use a test split for their dataset, but we hold back every $100$th image alphabetically for testing.
+We reevaluate H-3DGS for this scene with the new test split, using the camera calibrations and chunk splits provided on their website.
+In accordance with the instructions, we disable exposure optimization for evaluation.
+On this dataset, we use depth supervision identical to Hierarchical.
+Hierarchical 3DGS single-chunk dataset
+To demonstrate our ability to reconstruct small scenes, we also evaluate our method on the smaller, single-chunk versions of the Campus and Small City (cf.
+Table 4).
+The quality metrics indicate that training on the LoD structure (which would not be required for scenes of this scales) during training only leads to marginal reductions in visual quality.
+The distribution of street views in the campus dataset is visualized in Figure 20(b).
+Table 4: H-3DGS single chunk view synthesis results. Results reproduced from Hierarchical.
+Small City Chunk (1.1k images) Campus Chunk (1.5k images)
+[…]
+3DGS 25.34 0.776 0.337 23.87 0.785 0.378
+H-3DGS (leaves) 26.62 0.820 0.259 24.61 0.807 0.331
+Ours 25.94 0.808 0.276 24.86 0.796 0.360
+Mill19 dataset
+Table 5 shows results on the Mill19 dataset (MegaNerf).
+We use the camera poses and sparse point cloud provided by CityGaussian.
+In accordance with baselines, we downscale all images by a factor of 4.
+We have chosen this dataset, because it is widely used in large-scale novel-view synthesis.
+At the same time, this dataset represents a worst-case scenario for our method, as the regular, same-height aerial views (cf.
+Figure 20(a)) negate any benefit of our level-of-detail method and can be split into independent chunks trivially.
+Table 5: Mill19 novel view synthesis results Results with are reproduced from CityGaussian. Results of H-3DGS are reproduced from Hierarchical.
+Rubble (1.6k images) Building (1.9k images)
+[…]
+MegaNeRF 24.06 0.553 0.516 n/a 20.93 0.547 0.504 n/a
+3DGS 25.47 0.777 0.277 6.1M 20.46 0.720 0.305 6.4M
+CityGaussian 25.77 0.813 0.228 9.7M 21.55 0.778 0.246 13.2M
+H-3DGS 24.64 0.755 0.284 n/a 21.52 0.723 0.297 n/a
+Ours 23.75 0.704 0.330 25M 20.67 0.682 0.318 25M
+(a) rubble (Mill19)
+(b) campus (H-3DGS)
+Figure 20: Overview of the aerial and street datasets.
+A.4.1 Rendering Comparison
+To demonstrate the effectiveness of our LoD method, we compare VRAM usage and image quality to rendering the highest level of detail leaf Gaussians with both 3DGS (3DGS) and gsplat (Gsplat) using their default rasterization settings.
+We conducted comparisons across three different viewpoints to assess performance at varying levels of detail.
+As illustrated in Figure 21, our approach achieves visual fidelity comparable to the baselines while significantly reducing VRAM usage, highlighting the effectiveness of our LOD scheme.
+(a) Overview
+(b) Aerial View
+(c) Street View
+Figure 21: Comparision to standard 3DGS (3DGS) and GSplat (Gsplat) in terms of rendering quality and VRAM usage.
+A.5 Evaluation Details
+Finding baselines for the MC-small-city+ scene proved challenging, as the only methods we are aware of that have successfully trained on all fused aerial and street-level views–RetinaGS and NerfXL–both require 64 GPUs running in parallel.
+We were not able to meet these hardware requirements for evaluation and hold that these methods and similar multi-GPU works like Grendel are orthogonal to our method.
+H-3DGS (Hierarchical) has demonstrated the ability to reconstruct similar sized datasets, but its evaluation was restricted to street-level views.
+We choose Hierarchical and other popular large-scale Gaussian Splatting methods (CityGaussian; HorizonGS; OctreeGS) as baselines.
+Note that ”large-scale” is often used ambiguously in the context of novel-view synthesis, alternatively meaning reconstructing very complex scenes with large number of images or reconstructing novel views using far-away (i.e.
+aerial) input views: With the exception of Hierarchical, the biggest datasets these methods use in their papers consist of just 5620 images.
+In order to train MC-small-city+ with 42.2 thousand images, it was necessary to modify CityGaussian, HorizonGS and OctreeGS to load images from disk instead of caching them in RAM or VRAM.
+This slows down training, but does not affect the final results.
+Most large-scale 3DGS frameworks have complicated multi-stage training processes and are very sensitive to hyperparameters.
+We consider it a benefit of our training pipeline that the user is not confronted with tuning the error-prone partitioning process, being more akin to the original Gaussian Splatting training.
+In general, the hyperparameters suggested for large-scale scenes were used for baselines.
+If provided, we use configurations explicitly targeting the MatrixCity dataset, if it improves metrics over default configuration.
+Learning rates were not modified, as they were similar to ours and should not be a barrier to reconstruction on this scale.
+We also reached out to the authors of all our baselines about suggesting hyperparameters for the MC-small-city+ dataset, which we incorporated when provided.
+For methods that make use of configuration files, we include them in the supplemental material.
+OctreeGS
+For OctreeGS (OctreeGS), we follow their instructions on training custom datasets.
+Note that OctreeGS does not perform any scene division and uses the memory-intensive ScaffoldGS (ScaffoldGS) method, making it unsuitable for scenes of this scale.
+CityGaussian
+We choose CityGaussian (CityGaussian) (specifically the 1.2 version of the repository) as a baseline instead of CityGaussianV2 (CityGaussianV2), as the code release for CityGaussianV2 does not yet support level-of-detail rendering.
+We follow the instructions on training large datasets and use the parameters and chunk split from the provided configuration file for the MatrixCity-Aerial dataset (which covers the same region as the MC-small-city+ dataset, but without the street views), as they produced better metrics than the default parameters.
+Note that the low number of Gaussians in the final model is due to CityGS discarding most (about 90%) of the trained Gaussians after chunk training in order to avoid artifacts.
+For campus and small city, we used a $4\times 4$ and $2\times 2$ chunk split to match the number of chunks as closely as possible with the 1212/44 chunks used by H-3DGS (Hierarchical).
+Hierarchical-3DGS
+We follow the instructions of Hierarchical for running the method on large scenes.
+Note that we do not make use of their particular COLMAP pipeline for MC-smaller-city+, as it only works with unbroken camera paths.
+We disable exposure optimization for evaluation in accordance with the instructions.
+The partitioning step of H-3DGS requires point correspondences not provided by the dataset.
+To substitute, we use the camera poses provided by the dataset and generate the sparse point cloud and correspondences from scratch using COLMAP.
+On scenes of this scale, COLMAP output contains significant noise, which also leads to reduced quality metrics for our method.
+H-3DGS achieves drastically worse PSNR and SSIM results on the full campus dataset, compared to the single-chunk results, due to a slight perspective distortion that occurs specifically on this dataset (it did not occur on small city).
+This can be reproduced by evaluating the trained model on their campus dataset, both of which are available on their website.
+As such, we include the results, but encourage visual comparison in Figure 6.
+HorizonGS
+We expand the configuration provided for a single chunk of the MatrixCity small-scene dataset (named ours/large_scene/block_A) with a chunk split of $3\times 4$ and $5\times 5$ for MC-smaller-city+ and MC-small-city+ (shown in Figure 22) respectively.
+The partitions of the MC-small-city+ scene Each chunk is trained for 60000 iterations and then fine-tuned for another 40000.
+Figure 22: Partitions of HorizonGS on the MC-small-city+ scene.
+VastGaussian
+VastGaussian (VastGaussian) would have presented a natural comparison point, but there is currently no official code release available.
+Ours
+We provide detailed configuration files for all of the scenes as supplemental material.
+In general, we perform 60k iterations of coarse training (except for 100k on MC-small-city+) and then 150k iterations of fine training (except 250k on campus and MC-smaller-city+ and 500k on MC-small-city+).
+We shoose SH degree 1 and half the densifaction gradient threshold as opposed to H-3DGS.
+We pick a cache size of 15000000 Gaussians.
+A.6 Implementation Details
+Training models on this unprecedented scale on consumer hardware required solving many technical issues. In this section, we elaborate on the most important of these design decisions.
+Codebase
+Our general code structure is based on H-3DGS (Hierarchical).
+In particular, we reuse the hierarchy creator, but replace the cut procedure and chunk based training, as well as the rasterizer for the one used in 3DGS.
+We also use the densification implementation of 3DGSasMCMC.
+All of the code is implemented in PyTorch and C++/CUDA.
+The entire source code is included as supplemental material.
+Varying level of detail
+We introduce noise into the hierarchy cuts in order to prevent overfitting of Gaussians of a certain scale to views from a particular distance.
+In particular, we multiply the distances to the SPT centers each iteration with $1+5r^{4}$, with a uniformly random variable $r\in U(0,1)$.
+This factor is designed such that most iterations will train near the highest level of detail, but coarser levels will also be trained occasionally.
+While this does not improve image metrics on the test set, it shows significant improvement to out-of-distribution views and reduces LoD popping artifacts.
+Varying focal length
+The required level of detail is not only dependent on a camera’s distance $d$ to the Gaussian, but also on its focal length, which we need to account for as our datasets contain views with differing focal lengths.
+Therefore, we choose a base focal length $f_{b}$ and use a relative distance metric $\hat{d}$ for our HSPT cuts, which is calculated for the current camera’s focal length $f_{i}$ as $\hat{d}=\frac{f_{b}}{f_{i}}d$.
+E.g.
+for a camera with double the focal length, the cut distance should be halved to account for the larger projection footprint on the image plane.
+SH Degree
+For the most part, experiments indicate that SH degrees higher than 1 do not significantly contribute to image quality in the tested scenes.
+For training higher degrees $n$ of SH, we have found that increasing the degree from 0 to 1 during the course stage and then gradually increasing the degree from 1 to $n$ during fine training (with each increase happening after 10% of total training iterations) yields the best results.
+Gaussian Order
+To exploit spatial coherency and improve training performance, we store the Gaussians on CPU in Morton Z-order.
+As this order can change during training, we resort the Gaussians at every densification iteration.
+Training Images
+Conventional Gaussian Splatting stores the entire training dataset in VRAM to achieve their impressive training speed, however, this approach is infeasible for training ultra-large-scale datasets on consumer-grade hardware.
+Instead, we load the ground truth images from disk every iteration, saving VRAM at the cost of training performance.
+Unreachable Gaussians
+During training, it can occur that a child Gaussian becomes larger than its parent.
+This can lead to cases where $m_{d}(\text{parent}(i))<m_{d}(i)$, making it impossible for the parent Gaussian to fulfil the SPT cut condition.
+While we are aware of this inefficiency, we found that in practice this only happens to a small number of Gaussians ($10\%$).
+They still occupy a portion of RAM, but are never rendered or transferred.
+We have experimented with rebalancing the hierarchy during training, but found it too costly for little benefit.
+In general, we find this behaviour to be preferable over generating improper cuts, which can derail training.
+Respawning, Densifying, and Pruning Gaussians
+For improved performance, all respawn- and densify-operations are performed in parallel during the densification step.
+This can lead to difficult edge cases that need to be handled to prevent the hierarchy from degenerating: When two sibling nodes both need to be respawned, we only respawn the right node.
+This will cause the left node to become a new dead leaf, which will be respawned in the next densification iteration.
+If a node needs to be respawned whose sibling is not a leaf node, its entire subtree will replace the parent node.
+To minimize the number of unnecessary Gaussians, we apply a simple pruning strategy where we zero the opacity of Gaussians that have not contributed for a number of iterations equal to twice the size of the training set.
+This will cause them to be respawned in the next densification iteration.
+Performance Optimization
+The performance of the HSPT cut and Gaussian loading is particularly important, as they happen every iteration.
+We store the SPT properties ($m_{d}(i)$, $m_{d}(\text{parent}(i))$, $i$) for all Gaussians in a single continuous GPU memory buffer and perform the cuts in parallel using optimized CUDA kernels.
+Similarly, the Gaussian properties are stored in a single PyTorch tensor in RAM as an array of structures (each Gaussian’s properties concurrently).
+The properties of all required Gaussians are then transferred to the GPU via a single copy operation, reorganized on the GPU in a structure of arrays layout (as required by the rasterizer), and appended to the current render set.
+A.7 Additional Densification Details
+Our densification method combines the split operation from 3DGSasMCMC with the selection method from Hierarchical.
+Relying purely on the densification strategy of 3DGSasMCMC–applying a noise to every Gaussian’s position and an opacity/scaling loss–becomes problematic for large scenes, where a majority of Gaussians are not visible from any one view: The result is a gradual disappearance of density in the scene or extreme ”stringing” artifacts (cf.
+Figure 23).
+However, without these losses to encourage respawning of Gaussians, the densification becomes aimless and random.
+We implemented another strategy that only applies the losses and noise to Gaussians that affected at least one pixel in the current iteration.
+While this significantly improved results, we found that the MCMC densification strategy in general performs poorly when Gaussian density is low, which is necessitated by ultra-large-scale scenes.
+On small scenes such as the single-chunk datasets, our modified MCMC strategy slightly outperforms the presented strategy given sufficient Gaussian budget, but it lags behind in the larger scenes.
+(a) campus
+(b) small city
+Figure 23: Without regularization, the scale- and opacity-losses of 3DGS-as-MCMC lead to either stringing or disappearance of the entire scene.
+A.8 Pseudocode
+This section provides pseudocode for our core algorithms.
+Algorithm 1 contains a procedure to ”cut” a binary tree hierarchy along a certain condition.
+Algorithm 2 shows how to cut a single SPT at a particular distance $d^{j}$.
+The entire training procedure is sketched in Algorithm 3.
+Algorithm 1 BFS Hierarchy Cut
+procedure Cut(Hierarchy $\mathcal{H},condition$)
+  $cut\leftarrow\{\}$
+  $Q\leftarrow\textsc{Queue}(root(\mathcal{H}))$
+  while $Q$ not empty do
+   $node\leftarrow Q.\textsc{dequeue}()$
+   if $condition(node)$ then
+     $cut\leftarrow cut\cup node$
+   else
+     $Q.\textsc{enqueue}(node\rightarrow left)$
+     $Q.\textsc{enqueue}(node\rightarrow right)$
+   end if
+  end while
+  $\texttt{return }cut$
+end procedure
+Algorithm 2 SPT Cut
+procedure Cut(SPT $\mathcal{S}$, distance $d^{j}$)
+  $high\leftarrow\textsc{BinarySearch}(\mathcal{S}_{max},d^{j})$
+  $cut\leftarrow\{i\in[0,high]|\mathcal{S}_{min}^{i}<d^{j}\}$
+  $\texttt{return }cut$
+end procedure
+Algorithm 3 Full train procedure
+procedure Train(upper Hierarchy $\mathcal{U},cache,skybox$)
+  $view\leftarrow 0$
+  $Cache\_Distances\leftarrow\{\}$
+  $Cache\_SPTs\leftarrow\{\}$
+  while TrueTrue do
+   $view\leftarrow\textsc{Sample}(knn\_graph,view)$ $\triangleright$ Find a new nearby view
+   $cut\_condition\leftarrow[\lambda(i)=\neg\ is\_in\_frustum(\mu_{i})\lor m_{d}(i)<T]$
+   $upper\_cut\leftarrow\textsc{cut}(\mathcal{U},cut\_condition)$
+   $SPT\_roots\leftarrow\{u\in upper\_cut|u\text{ contains an SPT}\}$
+   $upper\_nodes\_to\_render\leftarrow upper\_cut\setminus SPT\_roots$
+   $SPT\_distances\leftarrow\{\left\|\mu_{r}-p_{view}\right\|_{2}^{2}\text{ for }r\in SPT\_roots\}$
+   $Cache\_Reuse\_SPTs\leftarrow\{s_{i}\in(Cache\_SPTs\cap SPT\_roots)|D_{min}\leq\left\|\frac{\mu_{i}-\mathbf{p}_{cam}}{Cache\_Distances_{i}}\right\|_{2}^{2}\leq D_{max}\}$
+   $load\_SPTs\leftarrow(SPT\_{roots}\setminus Cache\_Reuse\_SPTs)$
+   $load\_Gaussian\_indices\leftarrow\textsc{Cut\_SPTs}(load\_SPTs,SPT\_distances[load\_SPTs])$
+   $load\_Gaussians\leftarrow\textsc{Load}(load\_Gaussian\_indices)$
+   $Gaussians\leftarrow skybox\cup upper\_nodes\_to\_render\cup load\_Gaussians$
+   $Gaussians\leftarrow Gaussians\cup\textsc{load\_cache}(Cache\_Reuse\_SPTs)$
+   $\textsc{Render\_And\_Optimize}(Gaussians)$
+   $Cache\_SPTs\leftarrow Cache\_SPTs\cup load\_SPTs$
+   if len(Cache_Gaussians)>max_cache_sizelen(Cache\_Gaussians)>max\_cache\_size then
+     $\textsc{Write\_Back\_LRU}(Cache\_Gaussians)$ $\triangleright$ Write cached Gaussians to RAM
+   end if
+  end while
+end procedure
+A.9 Supplemental Material
+As supplemental material, we include the codebase used for this project, along with the configuration files for all of our scenes.
+We also include configuration files for our baselines, if applicable.
+Further, we provide videos at the following anonymous link: https://www.youtube.com/playlist?list=PLocO6QqSKZ4YN7DyuC5gWEeVciSV6RdAk.
+This includes fly-throughs of the evaluated scenes and a method overview video.
+A.10 Declaration of LLM usage
+LLMs were used to polish the language in the paper, but did not write any sections on their own. Some code sections in the implementation were generated using LLMs.

### v4

@@
-A LoD of Gaussians: Unified Training and Rendering for Ultra-Large Scale Reconstruction with External Memory
-Abstract
+A LoD of Gaussians: Out-of-Core Training and Rendering for Seamless Ultra-Large Scene Reconstruction
+Abstract.
 Gaussian Splatting has emerged as a high-performance technique for novel view synthesis, enabling real-time rendering and high-quality reconstruction of small scenes.
@@
 Together, these innovations enable seamless multi-scale reconstruction and interactive visualization of complex scenes—from broad aerial views to fine-grained ground-level details.
-1 Introduction
+Level of Detail, Gaussian Splatting, Large-Scale Reconstruction
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06submissionid: 395ccs: Computing methodologies Rasterizationccs: Computing methodologies Visibilityccs: Computing methodologies Machine learning
+Figure 1.
+Teaser: We introduce a fully hierarchical 3D Gaussian representation trained directly across unstructured, multi-scale image sets—including street-level and far aerial views—without scene partitioning.
+Our method maintains a consistent global scene model, eliminating boundary artifacts typical of chunked approaches.
+A hybrid Level-of-Detail system combines Gaussian hierarchies with Sequential Point Trees, enabling dynamic, view-dependent streaming and LoD selection.
+The entire model resides in external memory, with only a small, adaptive subset loaded on demand, allowing seamless training and interactive rendering of scenes with 150M+ Gaussians on a single consumer GPU ($\leq 24GB$ VRAM).
+1. Introduction
 Given a set of posed images of a 3D scene, the task of novel view synthesis (NVS) is to generate plausible images of the scene from unseen viewpoints.
-Early approaches achieved this via image-based blending (DeepBlending), but the introduction of Neural Radiance Fields (NeRF) (NeRF) marked a breakthrough, enabling high-quality results by optimizing an implicit volumetric scene representation through a multi-layer perceptron.
-More recently, 3D Gaussian Splatting (3DGS) (3DGS) extended this paradigm to an explicit representation: a set of Gaussian primitives that are efficiently rasterized using splatting techniques (zwicker2001EWA), replacing costly ray marching and allowing real-time rendering with fast convergence.
+Early approaches achieved this via image-based blending (Zhang et al., 2020), but the introduction of Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) marked a breakthrough, enabling high-quality results by optimizing an implicit volumetric scene representation through a multi-layer perceptron.
+More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) extended this paradigm to an explicit representation: a set of Gaussian primitives that are efficiently rasterized using splatting techniques (Zwicker et al., 2001), replacing costly ray marching and allowing real-time rendering with fast convergence.
 Despite these advances, both NeRF and 3DGS remain constrained by memory bottlenecks when applied to large-scale environments.
-Prior methods address this by dividing scenes into smaller chunks (GridNeRF; BlockNeRF; Hierarchical; CityGaussian; GigaGS; VastGaussian), training each independently before merging results.
+Prior methods address this by dividing scenes into smaller chunks (Xu et al., 2023; Tancik et al., 2022; Kerbl et al., 2024; Liu et al., 2024a; Chen et al., 2025; Lin et al., 2024), training each independently before merging results.
 While chunking strategies mitigate memory usage during training, they introduce several key limitations:
-  1. 1.
-    View-chunk misalignment: Camera views often span multiple chunks, especially in open environments or multi-scale datasets (e.g., combining aerial and street-level images).
-This makes chunk boundaries arbitrary and hard to define, complicating scene partitioning and training.
-We identify two artifacts caused by view-chunk misalignment: Chunk bleeding occurs when Gaussians extend out of their current chunk and obscure neighbouring chunks after merging.
-When an occluder on a training image is not part of the current chunk, its ”ghostly” outline will be trained into the wrong chunk, leading to Chunk ghosting.
-Examples of these artifacts can be seen in Figures6 and 11.
-  2. 2.
-    Redundant overlap: To avoid artifacts at chunk boundaries, regions are typically trained with significant overlap, which increases memory usage and prolongs training time.
-  3. 3.
-    Asymmetric hardware demands: Although chunking reduces memory requirements during training, rendering may require all visible chunks in memory simultaneously—often exceeding the capacity of the original training setup.
+  1. (1)
+    View-chunk misalignment: Camera views often span multiple chunks, especially in open or multi-scale datasets (e.g., combining aerial and street-level images).
+As a result, chunk boundaries become arbitrary with respect to images, complicating scene partitioning and training, resulting in artifacts: Chunk bleeding occurs when Gaussians extend out of their assigned chunk and obscure neighbouring chunks after merging.
+Chunk ghosting occurs when an occluder present in a training image is not part of the current chunk, training its ‘ghostly’ outline into the wrong chunk.
+See Figures 7 and 12.
+  2. (2)
+    Redundant overlap: To avoid artifacts at chunk boundaries, regions are typically trained with significant overlap, which duplicates parameters and optimizer state, increasing memory usage and prolonging training time.
+  3. (3)
+    Asymmetric hardware demands: Although chunking reduces memory requirements during training, rendering may require all visible chunks in memory simultaneously—often exceeding the capacity of the original training setup and undermining the practical benefit of partitioning.
 The simplest and most robust alternative to chunking is to avoid splitting altogether.
@@
 Still, a single distant view could require access to the full scene.
-To address this, we construct a hierarchical Level-of-Detail (LoD) model inspired by Hierarchical, loading detail proportional to view distance.
-Maintaining this hierarchy during training is non-trivial, as Gaussian properties evolve dynamically.
-We propose a novel hierarchy densification strategy, adapted from the MCMC-style spawning (3DGSasMCMC), to support stable, progressive refinement.
-Efficient view-dependent selection from the hierarchy is challenging for large models.
-Instead of full tree traversal, we adopt Sequential Point Trees (SPTs) (SPT), originally developed for point cloud LoD rendering.
-Our Hierarchical SPT version allows us to compute the correct LoD cut efficiently for rendering individual views and camera paths.
-Finally, to reduce CPU-GPU data transfer overhead, we introduce a lightweight caching system that tracks recently used Gaussians and reuses them across training iterations.
+To address this, we construct a hierarchical Level-of-Detail (LoD) model inspired by Kerbl et al.
+(2024), loading detail proportional to view distance.
+Crucially, this hierarchy must be maintained throughout training, as Gaussian parameters and their spatial distribution evolve dynamically.
+We propose a novel hierarchy densification strategy, inspired by MCMC-style spawning (Kheradmand et al., 2024), to support stable and progressive refinement.
+Efficient view-dependent selection from the hierarchy is challenging for large models, particularly when the hierarchy structure changes during optimization.
+Instead of full tree traversal, we adopt Sequential Point Trees (SPTs) (Dachsbacher et al., 2003), originally developed for point cloud rendering.
+Our Hierarchical SPT version allows us to compute the correct LoD cut efficiently for individual views and camera paths.
+Finally, to reduce CPU-GPU data transfer, we introduce a lightweight caching system that tracks recently used Gaussians and reuses them across training iterations.
 In summary, we make the following contributions:
-  1. 1.
-    We propose a novel hierarchy densification strategy that enables dynamic expansion and restructuring of Gaussian LoD representations during training.
-  2. 2.
-    We leverage the Sequential Point Tree (SPT) data structure and adapt it for large-scale Gaussian Splatting to perform fast, parallelizable LoD cuts for efficient training and rendering.
-  3. 3.
-    We demonstrate, for the first time, how external memory can be used to train and render ultra-large Gaussian scenes seamlessly on consumer-grade GPUs.
-  4. 4.
-    We design a caching and view scheduling system that exploits temporal coherence to minimize data transfer overhead and improve training throughput.
-  5. 5.
-    We verify trough extensive evaluations that the pipeline succeeds in reconstructing huge datasets with complex camera distribution, where SOTA divide-and-conquer methods fail even with larger VRAM budgets.
-2 Related Work
+  1. (1)
+    Seamless, non-partitioned training of ultra-large Gaussian scenes.
+We present the first 3D Gaussian Splatting framework that enables training and interactive rendering of city-scale scenes from arbitrary views without spatial partitioning, using out-of-core memory and view-dependent streaming on a single consumer GPU.
+  2. (2)
+    Dynamic LoD hierarchy densification during training.
+We propose a novel coarse-to-fine hierarchy densification strategy that allows Gaussian LoD hierarchies to evolve continuously during optimization, supporting stable refinement and restructuring without post-training hierarchy construction.
+  3. (3)
+    Hierarchical Sequential Point Trees for Gaussian Splatting.
+We adapt Sequential Point Trees (SPTs) to Gaussian splatting and introduce the Hierarchical SPT (HSPT), a hybrid data structure enabling efficient, parallelizable LoD selection while remaining robust to hierarchy updates during training.
+  4. (4)
+    Efficient out-of-core execution via cache-aware streaming and view scheduling.
+We design a lightweight GPU caching and view selection strategy that exploits temporal coherence to substantially reduce CPU–GPU transfer overhead during both training and rendering.
+  5. (5)
+    Large-scale evaluation and new benchmarks.
+We introduce the Uni10k dataset and an expanded version of the MatrixCity Small City scene, and demonstrate state-of-the-art performance on large-scale, multi-view datasets spanning aerial, street-level, and indoor environments.
+2. Related Work
 Large Scale Reconstruction
 Reconstructing large-scale scenes from images has long been a central challenge in visual computing.
-Traditional approaches relied on Structure-from-Motion (SfM) pipelines to recover geometry from unordered photo collections (RomeInADay; COLMAP).
-Differentiable rendering techniques, notably NeRF (NeRF) and 3DGS (3DGS), marked a paradigm shift by optimizing volumetric scene representations.
-Extensions of NeRF to large scenes typically employ scene partitioning (BlockNeRF; GridNeRF) or multi-GPU training strategies (NerfXL).
-Similarly, most large-scale 3DGS pipelines adopt chunk-based training: Hierarchical-3DGS (H-3DGS) (Hierarchical) trains chunks independently and then merges them into a global LoD hierarchy; CityGaussian (CityGaussian) combines chunked training with per-chunk LoD selection using LightGaussian (LightGaussian); and VastGaussian (VastGaussian) introduces decoupled appearance modeling and progressive partitioning.
-Horizon-GS (HorizonGS) integrates divide-and-conquer strategies with LoD mechanisms from OctreeGS, specifically targeting hybrid aerial/street-view datasets.
-GrendelGS (Grendel) avoids spatial chunking by distributing training images across GPUs, such that each device renders a disjoint screen region.
-Another research direction focuses on extracting geometric proxies from large-scale 3DGS scenes (ULSRGS; CityGaussianV2; GigaGS).
+Traditional approaches relied on Structure-from-Motion (SfM) pipelines to recover geometry from unordered photo collections (Agarwal et al., 2009; Schönberger and Frahm, 2016).
+Differentiable rendering techniques, notably NeRF (Mildenhall et al., 2021) and 3DGS (Kerbl et al., 2023), marked a paradigm shift by optimizing volumetric scene representations.
+Extensions of NeRF to large scenes typically employ scene partitioning (Tancik et al., 2022; Xu et al., 2023) or multi-GPU training strategies (Li et al., 2024b).
+Similarly, most large-scale 3DGS pipelines adopt chunk-based training: Hierarchical-3DGS (H-3DGS) (Kerbl et al., 2024) trains chunks independently and then merges them into a global LoD hierarchy; CityGaussian (Liu et al., 2024a) combines chunked training with per-chunk LoD selection using LightGaussian (Fan et al., 2024); OccluGaussian (Liu et al., 2025) partitions the scene to maximize camera correlation in each chunk; and VastGaussian (Lin et al., 2024) introduces decoupled appearance modeling and progressive partitioning.
+Horizon-GS (Jiang et al., 2025) integrates divide-and-conquer strategies with LoD mechanisms from Ren et al.
+(2024), specifically targeting hybrid aerial/street-view datasets.
+GrendelGS (Zhao et al., 2024) avoids spatial chunking by distributing training images across GPUs, such that each device renders a disjoint screen region.
+Another research direction focuses on extracting geometric proxies from large-scale 3DGS scenes (Li et al., 2025; Liu et al., 2024b; Chen et al., 2025).
 These methods leverage TSDF fusion and geometric losses to generate multiple meshes, which are fused and rendered efficiently using traditional rasterization.
+Other approaches (Zhao et al., 2026; Lee et al., 2025) make use out-of-core memory to scale training, but neglect LoD, which limits their memory reduction during training to frustum culling.
 Level-of-Detail Rendering
-Level-of-detail techniques reduce the geometric complexity of distant scene content to accelerate rendering.
-In the context of 3DGS, LoD approaches have been explored for enabling efficient rendering on memory-constrained or mobile devices.
-Compression-based strategies include attribute quantization via codebooks, pruning of low-impact Gaussians, and adapting the degree of spherical harmonics per primitive (ReducingTheMemoryFootprint; Compressed3DGaussianSplatting; LightGaussian; MiniSplatting; HierarchicalCompression; RadSplat; FLoD).
-Scaffold-GS (ScaffoldGS) uses latent vectors anchored to reference Gaussians, with an MLP generating associated Gaussians at render time.
-Octree-GS (OctreeGS) extends this concept to hierarchical LoD rendering, enabling real-time control over detail levels through spatial subdivision.
-3 Preliminaries
-Here, we briefly review 3DGS and how Gaussian hierarchies enable adaptive level-of-detail rendering.
-3.1 Hierarchical 3D Gaussian Splatting
-3DGS (3DGS) models a radiance field using a set of spatially distributed Gaussians, each with mean $\bm{\mu}_{i}\in\mathbb{R}^{3}$, RGB base colors $\mathbf{b}_{i}\in\mathbb{R}^{3}$ and covariance matrices $\bm{\Sigma}_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{\top}\mathbf{R}_{i}^{\top}$, which are parameterized via a diagonal scaling matrix $\mathbf{S}_{i}=\text{diag}(s^{1}_{i},s^{2}_{i},s^{3}_{i})$ and an orthonormal rotation matrix $\mathbf{R}_{i}$.
+Level-of-detail techniques reduce geometric complexity of distant scene content to accelerate rendering.
+In 3DGS, LoD approaches have mainly targeted efficient rendering on memory-constrained or mobile devices.
+Compression-based strategies include attribute quantization via codebooks, pruning low-impact Gaussians, and adapting the degree of spherical harmonics per primitive (Papantonakis et al., 2024; Niedermayr et al., 2024; Fan et al., 2024; Fang and Wang, 2024; Huang et al., 2025; Niemeyer et al., 2025; Seo et al., 2024).
+Scaffold-GS (Lu et al., 2024) uses latent vectors anchored to reference Gaussians, with an MLP generating associated Gaussians at render time.
+Octree-GS (Ren et al., 2024) extends this to hierarchical LoD rendering via spatial subdivision.
+Virtualized 3D Gaussians (Yang et al., 2025) targets composed scenes of reconstructed objects using an LoD scheme inspired by Unreal Engine 5’s Nanite (Karis et al., 2021).
+3. Preliminaries of Hierarchical 3D Gaussian Splatting
+3DGS (Kerbl et al., 2023) models a radiance field using a set of spatially distributed Gaussians, each with mean $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, RGB base colors $\mathbf{b}_{i}\in\mathbb{R}^{3}$ and covariance matrices $\boldsymbol{\Sigma}_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{\top}\mathbf{R}_{i}^{\top}$, which are parameterized via a diagonal scaling matrix $\mathbf{S}_{i}=\text{diag}(s^{1}_{i},s^{2}_{i},s^{3}_{i})$ and an orthonormal rotation matrix $\mathbf{R}_{i}$.
 Each Gaussian also stores an opacity $\sigma_{i}$ and view-dependent color, modeled using spherical harmonics (SH) coefficients $\mathbf{f}_{i}^{d}$.
@@
 For rendering, all $N$ Gaussians are sorted by distance to the camera and a discrete approximation of the volume rendering equation is evaluated for every pixel $\mathbf{x}$ with corresponding view direction $\mathbf{v}$:
-$\mathbf{C}(\mathbf{x})=\sum_{i=1}^{N}\mathbf{c}_{i}(\mathbf{v})\alpha_{i}(\mathbf{x})\prod_{j=1}^{i-1}(1-\alpha_{j}(\mathbf{x})),$ (1)
+(1) $\mathbf{C}(\mathbf{x})=\sum_{i=1}^{N}\mathbf{c}_{i}(\mathbf{v})\alpha_{i}(\mathbf{x})\prod_{j=1}^{i-1}(1-\alpha_{j}(\mathbf{x})),$
 where $\alpha_{j}$ is the opacity of the $j$-th Gaussian along the view ray:
-$\alpha_{j}(\mathbf{x})=\sigma_{j}e^{-\frac{1}{2}(\mathbf{x}-\bm{\mu^{\prime}}_{j})\bm{\Sigma^{\prime}}(\mathbf{x}-\bm{\mu^{\prime}}_{j})^{T}}.$ (2)
-Here $\bm{\mu^{\prime}}$ and $\bm{\Sigma^{\prime}}$ denote the projected 2D mean and covariance on the image plane, obtained by applying an affine approximation of the projective transform (zwicker2001EWA).
+(2) $\alpha_{j}(\mathbf{x})=\sigma_{j}e^{-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu^{\prime}}_{j})\boldsymbol{\Sigma^{\prime}}(\mathbf{x}-\boldsymbol{\mu^{\prime}}_{j})^{T}}.$
+Here $\boldsymbol{\mu^{\prime}}$ and $\boldsymbol{\Sigma^{\prime}}$ denote the projected 2D mean and covariance on the image plane, obtained by applying an affine approximation of the projective transform (Zwicker et al., 2001).
 3DGS Memory
-Standard 3DGS pipelines store the full set of per-Gaussian attributes, training images, and optimizer state in GPU memory (VRAM); see Figure A.1 for a detailed per-Gaussian breakdown.
+Standard 3DGS pipelines store the full set of per-Gaussian attributes, training images, and optimizer state in GPU memory (VRAM); see Figure 9 for a detailed per-Gaussian breakdown.
 Additional temporary allocations occur during forward and backward passes (e.g., for sorting and gradient accumulation).
@@
 Gaussian Hierarchies
-as introduced in H-3DGS (Hierarchical), recursively merge nearby Gaussians into a tree, where each non-leaf node approximates its children, and leaves correspond to the original Gaussians.
+As introduced in H-3DGS (Kerbl et al., 2024), Gaussian hierarchies recursively merge nearby Gaussians into a tree, where each non-leaf node approximates its children, and leaves correspond to the original Gaussians.
 A cut is defined by a condition $c_{\text{hier}}(i,\text{cam})$ evaluated in a breadth-first search (BFS).
-If a node satisfies the cut condition, it is added to the cut set and its children are skipped; otherwise, the BFS continues.
+If a node satisfies this condition, it is added to the cut set and its children are skipped; otherwise, the BFS continues.
 A proper cut set includes no parents or children of any included node and thus provides a view-adaptive LoD representation.
-The cut condition used in Hierarchical is a simple cutoff to the camera distance:
-$c_{\text{hier}}(i,\text{cam})=\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i),\quad m_{d}(i)=\frac{T}{\max_{j}s_{i}^{j}},$ (3)
+The cut condition used in (Kerbl et al., 2024) is a simple cut-off according camera distance:
+(3) $c_{\text{hier}}(i,\text{cam})=\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i),\quad m_{d}(i)=\frac{T}{\max_{j}s_{i}^{j}},$
 where $\mathbf{p}_{cam}$ is the camera position, $T$ is a global LoD threshold, and $m_{d}(i)$ is the minimum acceptable distance for viewing Gaussian $i$.
-The BFS ensures that if $i$ is in the cut set, parent($i$) failed the condition: $m_{d}(\text{parent}(i))>\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$
-4 Method
+The BFS ensures that if $i$ is in the cut set, parent($i$) failed the condition: $m_{d}(\text{parent}(i))>\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$
+4. Method
 To train models that exceed GPU memory limits, all Gaussian attributes are stored in CPU RAM and streamed to the GPU on demand for each training view according to the LoD hierarchy.
@@
 To minimize costly transfers between RAM and VRAM, we track which SPTs currently reside in GPU memory and at which detail.
-Only if an SPT is not present in this GPU cache at a similar level of detail, will it be loaded from RAM.
+Only if an SPT is not present in this GPU cache at a sufficiently similar level of detail, will it be loaded from RAM.
 Densification is performed on the CPU by adding new leaf nodes to the hierarchy and respawning low-opacity leaf nodes.
-The densified hierarchy is then converted back to an HSPT and transferred to the GPU for a new round of training iterations.
-Appendix A.2 includes more details on initialization and training.
-An overview of our training and densification process can be found in Figure 8 and pseudocodes in Appendix A.8.
-4.1 Sequential Point Trees for Gaussian Splatting
-Figure 1: An SPT and Gaussian hierarchy represent the same 5 Gaussians at varying levels of detail shown with four possible hierarchy cuts in red.
-Vertical lines in the SPT show the binary search result and horizontal lines the distance cut.
-Sequential Point Trees (SPT) were originally developed for point cloud LoD rendering, but can be trivially extended to ellipsoids and Gaussians.
-They enforce a more constrained but more efficient cut condition than Hierarchical:
-$c_{\text{SPT}}(i,\text{cam})=m_{d}(\text{parent}(i))>\left\|\bm{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$ (4)
-This condition is evaluated for all Gaussians in parallel, using the shared root-camera distance $\left\|\bm{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$.
+Following densification, the updated hierarchy is converted back to an HSPT and transferred to the GPU for a new round of training iterations.
+Suppl.
+A.2 includes more details on initialization and training.
+An overview of our training and densification process can be found in Figure 9 and pseudocode in Suppl.
+A.7.
+4.1. Sequential Point Trees for Gaussian Splatting
+(a) Hierarchy cuts
+(b) Densification example
+Figure 2.
+Visualization of Hierarchy and Densification.
+(2(a)) An SPT and Gaussian hierarchy show the same 5 Gaussians at varying levels of detail (red lines indicate cuts).
+Vertical lines show binary search results; horizontal lines show distance cuts.
+(2(b)) Leaf node densification and respawning.
+Sequential Point Trees (Dachsbacher et al., 2003) were originally developed for point cloud LoD rendering, but can be trivially extended to ellipsoids and Gaussians.
+They enforce a more constrained but more efficient cut condition than Kerbl et al.
+(2024):
+(4) $\scriptsize c_{\text{SPT}}(i,\text{cam})=m_{d}(\text{parent}(i))>\left\|\boldsymbol{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$
+This condition is evaluated for all Gaussians in parallel, using the shared root-camera distance $\left\|\boldsymbol{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$.
 It requires storing only sorted pairs $\big(m_{d}(i),m_{d}(\text{parent}(i))\big)$, significantly reducing memory compared to full Gaussian hierarchies.
@@
 Evidently, the level of detail of all Gaussians in the SPT is dictated by the camera’s distance to its root node.
-This can lead to Gaussians with $m_{d}(i)>\left\|\bm{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}$ being rendered, even though they would be too coarse for the current view.
-To counteract this issue, we define: $M_{d}(i)=m_{d}(i)+\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2},$ as a conservative minimum distance function.
-By the triangle inequality, selecting Gaussians satisfying $M_{d}(i)\leq\left\|\bm{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$ guarantees $m_{d}(i)\leq\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}$.
+This can lead to Gaussians with $m_{d}(i)>\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}$ being rendered, even though they would be too coarse for the current view.
+To counteract this issue, we define: $M_{d}(i)=m_{d}(i)+\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2},$ as a conservative minimum distance function.
+By the triangle inequality, selecting Gaussians satisfying $M_{d}(i)\leq\left\|\boldsymbol{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$ guarantees $m_{d}(i)\leq\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}$.
 In turn, this means that Gaussians that are further away from the camera than the root node will be selected at a higher level of detail.
@@
 Their compact memory footprint and parallel evaluation make them well-suited for large-scale scenes.
-Figure 1 visualizes both hierarchy types and their LoD cuts for a toy example.
-4.2 Densification
-Figure 2: Example of densifying and respawning leaf nodes.
+Figure 2 visualizes both hierarchy types and their LoD cuts.
+4.2. Densification
 Densifying an LoD representation presents a unique challenge, as the hierarchical structure must evolve continuously during training.
 Prior works circumvent this issue by constructing LoD hierarchies only after chunk-level training and densification are complete.
-We take inspiration from 3DGS-MCMC (3DGSasMCMC), which ‘splits’ Gaussians, replacing them with two new Gaussians which together should appear similarly to the original.
+We take inspiration from 3DGS-MCMC (Kheradmand et al., 2024), which ‘splits’ Gaussians, replacing them with two new Gaussians which together should appear similarly to the original.
 Notably, this mirrors how a parent node in a Gaussian hierarchy approximates its children.
-Therefore, we adopt this approach and instead ‘spawn’ two new child nodes for a Gaussian, increasing the size of the hierarchy with minimal artifacts.
-Instead of pruning, 3DGSasMCMC declares Gaussians below a certain opacity threshold as ‘dead’, and respawns them at the position of a high-opacity Gaussian.
+Motivated by this correspondence, we adopt this approach and instead ‘spawn’ two new child nodes for a Gaussian, increasing the size of the hierarchy with minimal artifacts.
+This procedure leads to gradual increase in detail during densification, thereby avoiding the instability on large scenes observed in (Kerbl et al., 2024).
+Instead of pruning, 3DGS-MCMC declares Gaussians below a certain opacity threshold as ‘dead’, and respawns them at the position of a high-opacity Gaussian.
 We propose a similar strategy: when a leaf node dies, its parent is replaced by its sibling node; the dead leaf node and its parent are then respawned as children to another node, which is selected to be densified.
-See Figure 2 for an overview of the two hierarchy densification operations.
-Together, they ensure that the hierarchy can be expanded during training in a stable manner and rebalanced as required.
-While 3DGSasMCMC choose Gaussians to densify using a random selection weighted by opacity, we employ the strategy from Hierarchical, which selects Gaussians for densification based on their maximal screen-space gradient.
-Avoiding destructive opacity resets, as advocated by 3DGS, is particularly important in our setting, as many Gaussians will no longer conform to the hierarchical structure after opacity is restored to normal.
-Further details on our densification method can be found in Appendix A.7.
-4.3 The hierarchical SPT datastructure
-This section covers previous approaches to LoD selection and why a new datastructure—the hierarchical SPT—is necessary for robust and efficient training.
+See Figure 2(b) for an overview of the two hierarchy densification operations.
+Together, they ensure that the hierarchy can be expanded during training in a stable and valid manner while being rebalanced as required.
+While 3DGS-MCMC choose Gaussians to densify using a random selection weighted by opacity, we employ the strategy from Kerbl et al.
+(2024), which selects Gaussians for densification based on their maximal screen-space gradient.
+This criterion better aligns densification with view-dependent reconstruction error in large-scale scenes.
+For further details on densification, see Suppl.
+A.6.
+4.3. The Hierarchical SPT Datastructure
+We first review previous LoD selection approaches and motivate the need for a new datastructure—the hierarchical SPT—for robust and efficient training.
 BFS
-Computing the cut set of a large Gaussian hierarchy is a costly operation that must be performed for every frame.
-A straightforward solution is to run a breadth-first search (BFS) from the root.
-This guarantees a proper cut and enables early pruning of large subtrees, e.g., via frustum culling.
-However, graph traversal is not well suited to parallel execution on the GPU.
+Computing the cut set of a large Gaussian hierarchy is costly and must be done for every frame.
+A straightforward solution is a BFS from the root, which guarantees a proper cut and enables early pruning of large subtrees (e.g., via frustum culling).
+However, graph traversal poorly suited to parallel execution on the GPU, making this approach prohibitively expensive at scale.
 Parallel Cut
-To enable GPU-accelerated cuts, Hierarchical evaluate the cut condition in parallel for each Gaussian:
-$\Big(m_{d}(i)<\left\|\bm{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}\Big)\land\Big(m_{d}(parent(i))\geq\left\|\bm{\mu}_{\text{parent}(i)}-\mathbf{p}_{cam}\right\|_{2}\Big),$ (5)
+To enable GPU-accelerated cuts, Kerbl et al. (2024) evaluate the cut condition in parallel for each Gaussian:
+$\Big(m_{d}(i)<\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}\Big)\land\Big(m_{d}(\text{parent}(i))\geq\left\|\boldsymbol{\mu}_{\text{parent}(i)}-\mathbf{p}_{cam}\right\|_{2}\Big)$,
 where any Gaussian that is sufficiently small at its current camera distance and whose parent is too large to be rendered, should be part of the cut set.
@@
 Hierarchical SPT (HSPT)
-Figure 3: A Gaussian hierarchy is converted to an HSPT by cutting according to Gaussian volume and converting sufficiently large subtrees to SPTs. The HSPT can then be cut in a 2-step process.
-We present the hierarchical SPT data structure, which combines the benefits of both approaches.
+Figure 3. A Gaussian hierarchy is converted to an HSPT by cutting according to Gaussian volume and converting sufficiently large subtrees to SPTs. The HSPT can then be cut in a 2-step process.
+Our HSPT data structure combines the benefits of both approaches.
 To construct it from a Gaussian hierarchy, we cut it using a BFS on the condition $c_{\text{HSPT}}(i)=s_{i}^{1}\cdot s_{i}^{2}\cdot s_{i}^{3}<\texttt{size}$ with volume threshold size.
@@
 This provides an upper bound on the error introduced if the subtree is converted into an SPT.
-Consequently, each subtree of sufficient size in $\mathbb{C}_{\text{HSPT}}$ can be transformed into a Sequential Point Tree to accelerate cut computation.
+Consequently, each subtree of sufficient size in $\mathbb{C}_{\text{HSPT}}$ can be transformed into an SPT to accelerate cut computation without violating cut correctness at higher levels.
 The HSPT-based cutting process then proceeds in two steps: first, a BFS on the upper hierarchy selects the required nodes and leaf/SPT subtrees for the current view.
@@
 The construction and cutting process of an HSPT is illustrated in Figure 3.
-Figure 10 and Figure 15 show example frames with respectively SPTs highlighted and different levels of detail.
-Rebuilding the HSPT for every training iteration would eliminate any performance benefit.
-Instead, we exploit the fact that the minimum distance $m_{d}$ evolves slowly during optimization and thus only needs to be updated infrequently.
+Figure 11 and 14 show example frames with respectively SPTs highlighted and different levels of detail.
+Rebuilding the HSPT every training iteration would eliminate its performance benefits.
+Instead, we exploit the fact that the minimum distance $m_{d}$ evolves slowly during optimization and thus requires only infrequent updates.
 In practice, we rebuild the HSPT only after each densification step.
-This infrequent recomputation allows us to use a more accurate—albeit more expensive—minimum distance metric than the inverse of maximal scale. Specifically, we define:
-$m_{d}^{\prime}(i)=\frac{T}{\sqrt{s_{i}^{1}\cdot s_{i}^{2}+s_{i}^{1}\cdot s_{i}^{3}+s_{i}^{2}\cdot s_{i}^{3}}},$ (6)
+This infrequent recomputation allows us to use a more accurate—albeit more expensive—minimum distance metric than the inverse of maximal scale.
+Specifically, we define:
+(6) $m_{d}^{\prime}(i)=\frac{T}{\sqrt{s_{i}^{1}\cdot s_{i}^{2}+s_{i}^{1}\cdot s_{i}^{3}+s_{i}^{2}\cdot s_{i}^{3}}},$
 which corresponds to the inverse square root of the surface area of the Gaussian ellipsoid (up to a constant factor).
@@
 Frustum Culling
-The major benefits of using BFS to cut the upper hierarchy are the guarantee of a proper cut and early culling of subtrees.
-To this end, we frustum cull every node considered in the BFS by checking if a sphere around the Gaussians with radius equal to $(3\cdot\max_{j}s_{i}^{j})$ intersects with the view frustum.
-Using the Gaussian scale as a proxy for the extent of its entire subtree is not perfectly accurate, but comparisons to using a full bounding sphere hierarchy showed no discernable difference in our experiments.
-It should be noted that Gaussians are implicitly frustum culled during rasterization, but this early culling accelerates the cutting procedure and significantly decreases the number of Gaussians that need to be loaded from RAM, as can be seen in Figure 9.
-4.4 Caching on the GPU
-Figure 4: Gaussians required for the current training view are assembled from three sources: the upper tree, newly loaded SPT cuts from RAM, and cache hits.
-After optimization, newly accessed SPTs are added to the GPU cache.
+The main benefits of using BFS to cut the upper hierarchy are the guarantee of a proper cut and early culling of subtrees.
+We therefore frustum cull every node considered in the BFS by checking if a sphere around the Gaussians with radius $3\cdot\max_{j}s_{i}^{j}$ intersects the view frustum.
+While using the Gaussian scale as a conservative proxy for the entire subtree extent is not perfectly accurate, we observed no discernable difference compared to a full bounding sphere hierarchy in our experiments.
+Although Gaussians are implicitly frustum culled during rasterization, this early culling accelerates cut computation and significantly reduces the number of Gaussians loaded from RAM (see Figure 10).
+4.4. Caching on the GPU
 Loading Gaussian data from RAM is a costly operation that can become a significant bottleneck during large-scale training.
@@
 Rather than caching individual Gaussians, we store the Gaussians from SPT cuts along with the cached distance from the camera to the root of each SPT, denoted $\bar{d}^{j}$ for the $j$-th SPT.
-During rendering, when the upper hierarchy is cut and the required SPTs identified, we compute $d^{j}=\|\bm{\mu}_{\text{root}(j)}-\mathbf{p}_{\text{cam}}\|_{2}$ and check whether a matching cut is cached, using a simple distance ratio tolerance:
-$D_{\text{min}}\leq\frac{d^{j}}{\bar{d}^{j}}\leq D_{\text{max}}.$ (7)
+During rendering, when the upper hierarchy is cut and the required SPTs identified, we compute $d^{j}=\|\boldsymbol{\mu}_{\text{root}(j)}-\mathbf{p}_{\text{cam}}\|_{2}$ and check whether a matching cut is cached, using a simple distance ratio tolerance:
+Figure 4.
+Gaussians required for the current training view are assembled from three sources: the upper tree, newly loaded SPT cuts from RAM, and cache hits.
+After optimization, newly accessed SPTs are added to the GPU cache.
+(7) $D_{\text{min}}\leq\frac{d^{j}}{\bar{d}^{j}}\leq D_{\text{max}}.$
 Here, $D_{\text{max}}$ defines the allowable range for using coarser-than-optimal detail, while $D_{\text{min}}$ limits how much finer detail can be tolerated.
@@
 While this heuristic introduces slight variability in rendered detail—since the LoD may depend on the cache state—we find that this stochasticity actually improves training robustness.
-In particular, the subtle variation in detail across views discourages overfitting to fixed camera distances and promotes better generalization across scales.
+In particular, subtle variations in detail across views discourage overfitting to fixed camera distances and promotes generalization across scales.
 For each training view, visible Gaussians are assembled from the upper hierarchy, the cached SPTs, and the skybox Gaussians (which remains in VRAM).
@@
 The next training view $j$ is then sampled from the $k$-nearest neighbours of the current view $i$ according to the distribution: $\mathbb{P}(j\mid i)\propto\frac{1}{w_{ij}+W},$ where $W$ is a normalization constant that also controls the degree of exploration.
-However, care must be taken when modifying the order of training views, as deviating from uniform random sampling may introduce training bias.
+However, care must be taken when deviating from a uniformly sampled training view selection as this may introduce bias.
 To counteract this, we inject a randomly selected view every 128 iterations, which we find sufficient to preserve generalization performance.
-4.5 Memory Layout
-Figure 5: Peak memory consumption of CPU and GPU for a training iteration on MC-smaller-city+ with 60 million Gaussians.
+4.5. Memory Layout
+Figure 5. Peak memory consumption of CPU and GPU for a training iteration on MC-smaller-city+ with 60 million Gaussians.
 Figure 5 illustrates the peak memory usage for a single training iteration of a 60-million-Gaussian hierarchy on the MC-smaller-city+ dataset.
-The output frame from this iteration is shown in Figure 13.
 The majority of RAM usage is consumed by per-Gaussian properties and their corresponding ADAM optimizer states.
@@
 The remaining GPU memory usage consists of auxiliary data, including cache management, hierarchy cut tracking, training and ground-truth images, as well as general PyTorch overhead.
-Figure A.1 provides a breakdown of all data associated with a single Gaussian and how they are distributed across CPU and GPU.
-5 Evaluation
+Figure 9 provides a breakdown of all data associated with a single Gaussian and how they are distributed across CPU and GPU.
+Table 1.
+Novel view synthesis results.
+Results with require suboptimal COLMAP initialization instead of the provided point cloud.
+VRAM usage is measured while rendering the test images.
+Methods are considered out-of-memory (OOM) if they exceed 141GB of VRAM during training or rendering.
+[…]
+MC-small-city+ (42.2k images)
+CityGaussian 19.78 0.650 0.475 2.4GB 4.1GB 1.11M 2.7M
+HorizonGS 12.06 0.521 0.544 20.6GB 25.3GB 2.5M N/A
+OctreeGS OOM OOM OOM OOM OOM OOM OOM
+H-3DGS OOM OOM OOM OOM OOM OOM OOM
+CLM 15.68 0.584 0.523 18.7GB 25.1GB 600k 24.2M
+Ours 21.59 0.711 0.396 16.4GB 18.0GB 600k 136.7M
+MC-smaller-city+ (15.1k images)
+HorizonGS 15.28 0.601 0.435 13.9GB 23.6GB 1.2M 34.1M
+OctreeGS 15.93 0.560 0.556 26.6GB 98.7GB 310k N/A
+Ours 21.71 0.710 0.370 15.7GB 17.2GB 310k 60M
+H-3DGS 14.57 0.516 0.563 97.0GB 15.1GB 4M 83.1M
+Ours 20.59 0.658 0.468 20.7GB 22.1GB 310k 24.7M
+Uni10k (10.1k images)
+CityGaussian 19.78 0.642 0.414 3.7GB 14.6GB 300k 2.2M
+H-3DGS 19.96 0.705 0.319 39.2GB 26.5GB 255k 32.5M
+OctreeGS 20.06 0.602 0.341 25.3GB 22.9GB 160k N/A
+CLM-GS 19.89 0.608 0.455 46.7GB 39.7GB 160k 12.6M
+HorizonGS 21.72 0.722 0.326 10.3GB 44.3GB 400k N/A
+Ours 21.32 0.718 0.336 22.6GB 31.3GB 160k 28.6M
+5. Evaluation
 Our method is designed to enable seamless training and rendering on ultra-large-scale scenes comprising tens of thousands of views captured at vastly different scales.
 Unfortunately, most datasets of sufficient size contain either street-level or aerial views—but not both.
-To overcome these challenges, we turn to the synthetic MatrixCity dataset (MatrixCityDataset), which provides dense multi-scale views along with ground-truth camera poses and sparse geometry.
-We also include results on the real-world street-level datasets from H-3DGS.
-As baseline, we compare to recent large-scale 3DGS methods CityGaussian (CityGaussian), H-3DGS (Hierarchical), HorizonGS (HorizonGS) and OctreeGS (OctreeGS).
-We note that HorizonGS is not evaluated on the H-3DGS dataset, since it requires aerial views for training.
-Further results and additional details on the datasets and baselines and are included in A.4.
-Evaluations were run on a single H200 GPU with 141 GB VRAM, while all of our models were trained on a single RTX 3090 graphics card with 24 GB VRAM.
-The contribution of our components is ablated in Appendix A.3.
-5.1 Results
+To address this gap, we captured the campus of Udine University with over 10 000 images at 4k resolution from varying aerial heights and street views in the Uni10k scene.
+We also introduce MC-small-city+, which expands the small-city scene from the MatrixCity (Li et al., 2023) dataset with new high-altitude aerial views covering the entire scene, resulting in 42.2k images spanning hundreds of buildings.
+Since some methods run out-of-memory on scenes of this scale, we additionally construct a subset of MC-small-city+ covering about a third of the area (15.1k images, MC-smaller-city+).
+Additionally, we present results on large-scale street-view and indoor scenes from the Hierarchical 3DGS (Kerbl et al., 2024) and OccluGaussian (Liu et al., 2025) datasets.
+Results on the aerial datasets UrbanScene3D (Lin et al., 2022) and Mill19 (Turki et al., 2022), as well as the smaller-scale H-3DGS single chunk scene are included in Suppl.
+A.3 along with additional details on all scenes.
+We choose recent divide-and-conquer based 3DGS methods CityGaussian (Liu et al., 2024a), H-3DGS (Kerbl et al., 2024), the large-scale neural Gaussian method OctreeGS (Ren et al., 2024) and out-of-core training method CLM-GS (Zhao et al., 2026) as baselines.
+Because OccluGaussian (Liu et al., 2025) has not released code yet, we compare only against their self-reported results on their dataset.
+To enable training on scenes of this scale, it was necessary to modify CityGaussian, HorizonGS and OctreeGS to load images from disk instead of caching them fully in RAM or VRAM.
+These changes affect training throughput but do not alter optimization behavior or final reconstruction quality.
+In general, the hyperparameters suggested for large-scale scenes were used for the experiments.
+Exact details on the training setting and configuration files are included in the supplemental material.
+Evaluations were run on a single H200 GPU with 141 GB VRAM to allow baselines methods to complete training.
+Where applicable, we report rendering performance on consumer GPUs to demonstrate practical deployability.
+5.1. Results
 Qualitative comparisons are shown in Figure 12 and quantitative comparisons in Tables 1 and 2.
-The outputs of H-3DGS and HorizonGS show significant floating and ghosting artifacts.
+The outputs of H-3DGS show significant floating and ghosting artifacts.
 While H-3DGS performs well on individual chunks of MC-smaller-city+, oversized floaters survive the merging procedure and obscure the majority of test images (cf.
-Figure 11, 13 and 17).
-The merging procedure of CityGaussian (CityGaussian), which is designed for aerial-only datasets, discards most of the trained Gaussians to avoid chunk artifacts, leading to mostly artifact-free views, but significantly blurry results.
-A LoD of Gaussians reconstructs novel views with a high degree of detail while avoiding chunk-based artifacts (cf.
-Figure 16), resulting in a considerable lead in quality metrics.
-Seamless training enables faster convergence during training, reducing the required number of training iterations considerably over divide-and-conquer based methods.
-Table 1: MatrixCity novel view synthesis results.
-Results with * use the MC-smaller-city+ dataset due to memory constraints.
-Results with require suboptimal COLMAP initialization instead of the dataset’s provided point cloud.
-VRAM usage is measured while rendering the test images.
-Methods are considered out-of-memory (OOM) if they exceed 141GB of VRAM during rendering or training.
-MC-small-city+ (42.2k images) / *MC-smaller-city+ (15.1k images)
+Figure 7).
+Floaters and ghosting artifacts can also be found in the results on other scenes.
+The merging procedure of CityGaussian, which is designed for aerial-only datasets, discards most of the trained Gaussians to avoid chunk artifacts, resulting in mostly artifact-free views and low memory requirements, but noticeably blurry results.
+A LoD of Gaussians reconstructs novel views for joined street- and aerial data with a high degree of detail while avoiding chunk-based artifacts, resulting in a consistent improvement in quality metrics across scenes.
+Seamless training also enables faster convergence, substantially reducing the number of training iterations required compared to divide-and-conquer based methods.
+Results on the OccluGaussian dataset further demonstrate generalizability to indoor environments, which are particularly difficult for divide-and-conquer based methods.
+All experiments of our method were run on an RTX 3090 GPU, except Uni10k, which required more VRAM for the 4k training images.
+Lowering the resolution to HD reduces peak VRAM to 20GB for the same parameters.
+CLM-GS also employs out-of-core memory during training, but relies solely on frustum culling to limit memory pressure.
+On Uni10k and MC-small-city+, this results in substantially higher VRAM usage per Gaussian than our method.
+Further, unlike OctreeGS (Ren et al., 2024) and our approach, their densification is not tailored to large-scale scene reconstruction from sparse views and point clouds.
+On MatrixCity, CLM-GS instead initializes from a fully dense synthetic point cloud; larger non-aerial scenes (e.g.
+Campus) fail to reconstruct under all configurations provided with their code.
+HorizonGS (Jiang et al., 2025) is not competitive on the MC-city+ tests due to the large scale of the dataset and its diversity of viewpoints.
+While the authors report results for MatrixCity, only a single block is used in their evaluation.
+However, it performs competitively on Uni10k, for which it is explicitly designed: the method targets mixed street- and aerial-view scenarios and relies on ground-truth labels to primarily target street-level views for densification.
+Our universal approach—without additional supervision—closely matches HorizonGS on Uni10k and drastically outperforms it on both MC-city+ scenes.
+Furthermore, rendering the neural HorizonGS representation is impractical due to a costly preprocessing step: while enabling low VRAM usage, it results in render times of up to 10 seconds per frame.
 […]
-CityGaussian 19.78 0.650 0.475 2.4 GB 1.11M 2.7M
-HorizonGS 12.06 0.521 0.544 20.6 GB 2.5M 50.3M
-H-3DGS OOM OOM OOM >>141GB OOM OOM
-OctreeGS OOM OOM OOM >>141GB OOM OOM
-Ours 21.59 0.711 0.396 25.0GB 600k 136.7M
-HorizonGS* 15.28 0.601 0.435 13.91 1.2M 34.1M
-OctreeGS* OOM OOM OOM >>141GB OOM OOM
-Ours* 21.71 0.710 0.370 17.7GB 310k 60M
-H-3DGS* 14.57 0.516 0.563 97GB 4M 83.1M
-Ours* 20.59 0.658 0.468 21.7GB 310k 24.7M
-Table 2: Hierarchical 3DGS dataset novel view synthesis results.
-Small City (5.8k images) Campus (22.0k images)
+Dataset Hierarchical 3DGS OccluGaussian
+Scene Small City (5.8k) Campus (22.0k) Canteen (8.6k) Classbuilding (9.1k)
 […]
-OctreeGS 19.20 0.601 0.498 OOM OOM OOM
-CityGS 21.40 0.692 0.383 19.47 0.645 0.531
-H-3DGS (leaves) 24.35 0.788 0.273 17.84 0.6101 0.4667
-Ours 24.61 0.774 0.297 21.85 0.707 0.456
-(a) Ground Truth
-(b) Ours
-(c) CityGaussian
-(d) Hierarch. 3DGS
-(e) HorizonGS
-Figure 6: Qualitative comparison of our method and SOTA methods (CityGaussian; Hierarchical; OctreeGS) on the MC-small-city+ and campus datasets.
+OctreeGS 19.20 0.601 0.498 20.54 0.643 0.544 21.49 0.826 0.293 23.08 0.889 0.197
+CityGS 21.40 0.692 0.383 19.47 0.645 0.531 20.41 0.794 0.275 20.48 0.840 0.244
+H-3DGS 24.35 0.788 0.273 17.84 0.610 0.467 22.71 0.825 0.178 23.87 0.881 0.128
+CLM-GS 22.54 0.699 0.392 16.60 0.613 0.558 23.13 0.864 0.168 22.89 0.892 0.159
+OccluGS N/A N/A N/A N/A N/A N/A 25.25 0.900 0.100 25.33 0.921 0.083
+Ours 24.61 0.774 0.297 21.85 0.707 0.456 23.33 0.867 0.155 24.14 0.915 0.116
 Rendering
-Our level-of-detail and caching strategy can also be applied to efficiently render the finished models: As supplemental material, we include fly-through videos of all scenes rendered on an RTX 3090 GPU at this anonymized link: https://www.youtube.com/playlist?list=PLocO6QqSKZ4YN7DyuC5gWEeVciSV6RdAk.
-Table 1 demonstrates effective VRAM reduction during rendering of test images in comparison with other large-scale 3DGS methods.
-Comparisons of rendering VRAM with 3DGS (3DGS) and gsplat (Gsplat) are included in Appendix A.4.1.
-6 Discussion and Outlook
-A LoD of Gaussians enables seamless training and rendering of ultra-large 3DGS models on consumer hardware.
-By storing Gaussian data in external memory and streaming it on demand, our method avoids the pitfalls of chunk-based pipelines.
-The hierarchical SPT structure accelerates LoD selection and remains robust to ongoing training changes.
-Combined with caching and view selection, our system significantly reduces out-of-core overhead.
-Together, these components enable efficient reconstruction and rendering at scale, demonstrated on the MC-small-city+ dataset.
-Limitations and Future Work
-A primary limitation of our method lies in initialization.
-Accurate estimation of camera poses and sparse point clouds remains challenging at large scale, particularly for real-world datasets with sparse or inconsistent coverage.
-Future work could explore improved initialization strategies or joint optimization of poses and Gaussians during early training.
-Our system also requires roughly 1 GB of RAM per million Gaussians, which—while more efficient than prior methods—still constrains scalability.
-Loading from disk is feasible in our experiments, though at the cost of a $10\times$ slowdown.
-Further, while our method in general greatly reduces the number of training iterations necessary, individual training iterations can take a lot longer than standard 3DGS due to the loading overhead.
-The level-of-detail system makes our method robust against vast differences in view distance, but when these differences are not present, like in single-height aerial datasets as evaluated in Appendix A.4, it becomes a burden instead, making more straight-forward divide-and-conquer training the better choice.
-Interactive rendering could be further optimized by avoiding per-frame hierarchy cut recomputation and enabling asynchronous streaming.
-While frustum culling reduces memory load in most views, it is ineffective when the entire scene falls inside the frustum.
-Occlusion culling could address this by skipping entire SPTs before loading.
-Overall, we believe that out-of-core 3D Gaussian Splatting is a promising direction for scaling radiance field methods to city-scale and beyond, without the need for specialized hardware.
-Reproducibility Statement
-To ensure reproducibility of our results, we have appended our entire codebase along with the config files of our method and baselines required to train each of the scenes.
-Detailed breakdowns of datasets and hyperparameters can be found in Appendix A.4 and A.5.
-Appendix A Appendix
-A.1 Figures
-(a) Divide and Conquer
-(b) Ours
-Figure 7: Comparison between the divide and conquer training process as used in Hierarchical; VastGaussian; CityGaussian; HorizonGS to our training process.
-Colored regions are present in VRAM, training views are drawn in red.
-[…]
-Figure 9: Frustum Culling and LoD selection (left) greatly reduces the number of Gaussians required to render a view (right).
-Figure 10: Hierarchical SPTs enable smooth transitions between detailed (left) and coarse (right) representation.
-(a) Chunk Bleeding in H-3DGS (Hierarchical).
-(b) Chunk Ghosting in CityGaussian (CityGaussian).
-Figure 11: Artifacts caused by the divide-and-conquer strategy on the MC-small-city+ dataset.
-Figure 12: Training image rendered during the iteration depicted in Figure 5.
-Figure 13: Images from H-3DGS on the MatrixCity-Scale scene are obscured by large floaters.
-Figure 14: All Gaussian properties except gradients are always present in CPU RAM, whereas only the currently loaded Gaussians and slim SPT information needs to be stored on GPU.
-Figure 15: SPTs for a frame of MatrixCity rendered in different colors.
-Figure 16: A LoD of Gaussians is able to reconstruct and render huge environments with 150M+ Gaussians with large variations of scale on consumer hardware.
-(a) Image generated during chunk training by H-3DGS.
-(b) The same image generated during training by A LoD of Gaussians.
-Figure 17: Chunk-based optimizations causes floaters to expand outside of the chunk boundary.
-A.2 Initialization and Training
-Following Hierarchical, we initialize the Gaussian model from a sparse point cloud, augmented with skybox points.
-This initial representation is small enough to fully reside in GPU memory and is trained for 100100k iterations without densification.
-The goal of this phase is to establish a stable global scene structure before constructing the hierarchy.
-After this initial optimization, we build a binary Gaussian hierarchy, where the trained Gaussians act as leaf nodes and the parent nodes represent merged approximations of their two children.
-The Gaussian properties–including base color, SH coefficients, position, and covariance–are optimized using the standard loss propagation introduced by 3DGS.
-To fulfill our tight memory requirements, we update only the parameters of the cut set chosen for each training view.
-Thus, gradients may, in general, be propagated to Gaussians in the middle of the hierarchy.
-Assuming good coverage and diversity of training views, these updates will diffuse over the entire hierarchy over the course of training, leading to smooth transitions from the highest to the lowest LoD (see Figure 10).
-A.3 Ablations
-We assess the contribution of key components through an ablation study, using recorded camera paths across all scenes (see supplemental videos).
+Our level-of-detail and caching strategy can also be applied to efficiently render the trained models.
+As supplemental material, we include fly-through videos of the evaluated scenes rendered on an RTX 3090.
+Table 1 demonstrates effective VRAM reduction during rendering of test images compared to other 3DGS methods.
+Figure A LoD of Gaussians: Out-of-Core Training and Rendering for Seamless Ultra-Large Scene Reconstruction compares rendering VRAM consumption and image quality of our LoD method with full-detail 3DGS (Kerbl et al., 2023) and gsplat (Ye et al., 2025).
+Our approach achieves visual fidelity comparable to the baselines while significantly reducing VRAM usage, highlighting the effectiveness of our LoD scheme.
+Ablations
+We assess the contribution of key components through an ablation study using recorded camera paths across a selection of scenes (see supplemental videos).
 Table 3 reports average frame times over these paths.
-Caching significantly improves performance, roughly doubling the framerate across scenes.
-Frustum culling is particularly beneficial for large-scale models but may incur overhead in smaller scenes.
+Caching significantly improves rendering performance, roughly doubling the framerate across scenes by reducing the average number of Gaussians loaded from RAM by 93% on Campus and 86% on MC-small-city+.
+The effectiveness of frustum culling scales with scene size: On the full MC-small-city+ scene, 24.5 million Gaussians are frustum culled on average (88% reduction), while for Campus and Small City the corresponding values are 9.8 million and and 7.9 million (74% and 65% reduction), respectively.
+As expected, the overhead of frustum culling amortizes with scene size, making it essential for the largest datasets.
 To evaluate cut efficiency, we compare the time required to compute the visible set using either full hierarchy BFS or our HSPT-based approach.
-HSPT consistently yields faster cut times.
-For training, we measure average iteration durations over 1 000 steps.
-Here, frustum culling and caching prove essential, substantially reducing the number of Gaussians loaded and processed per view.
-Table 3: Ablations.
+HSPT consistently yields faster cut times due to improved parallelization.
+Moreover, the BFS approach requires positions and scales for all Gaussians to reside in memory, causing it to exceed 24GB of VRAM on MC-small-city+, whereas the HSPT method peaks at 21GB.
+For training ablations, we measure average iteration durations over 1 000 steps.
+Here, frustum culling and caching prove essential, substantially reducing the number of Gaussians loaded and rendered per view.
+Table 3.
+Ablations.
 Average frame times for rendering camera paths and average iteration times during training with and without caching Gaussians.
 The final results show the average timings of the hierarchy cut during rendering using our HSPT and the baseline BFS approach.
-For campus, we evaluate two differently sized models with 38M and 80M Gaussians respectively.
+For Campus, we evaluate two different models with 38M and 80M Gaussians respectively.
 MC-smaller-city+ MC-small-city+ Campus
@@
 Render w/o Cache 119.4 ms 127.4 ms 92.6 ms 222.3 ms
-Render w/o Frustum Culling 52.3 ms 89.3 38.3 ms 110.2 ms
-Train Full 156 ms 174 ms 205 ms
-Train w/o Cache 471 ms 257 ms 244 ms
-Train w/o Frustum Culling 685 ms 626 ms 312 ms
-HSPT Cut (Render) 31.9 ms 45.7ms 31.3 ms 36.5 ms
-BFS Cut (Render) 47.8 ms OOM 40.0 ms 53.7 ms
+Render w/o Frustum Culling 52.3 ms 89.3 ms 38.3 ms 110.2 ms
+Training Full 156 ms 174 ms 205 ms
+Training w/o Cache 471 ms 257 ms 244 ms
+Training w/o Frustum Culling 685 ms 626 ms 312 ms
+Render Cut (HSPT) 31.9 ms 45.7ms 31.3 ms 36.5 ms
+Render Cut (BFS) 47.8 ms OOM 40.0 ms 53.7 ms
+6. Discussion and Outlook
+A LoD of Gaussians enables seamless training and rendering of ultra-large 3DGS models on consumer hardware.
+By storing Gaussian data in external memory and streaming it on demand, our method avoids the pitfalls of chunk-based pipelines.
+The HSPT datastructure accelerates LoD selection and remains robust to ongoing training changes.
+Combined with caching and view selection, our approach significantly reduces out-of-core overhead.
+These components enable efficient reconstruction and rendering at scale, as demonstrated on challenging multi-scale scenes such as MC-small-city+.
+Limitations and Future Work
+Our method represents an informed trade-off between performance and memory.
+While it greatly reduces the necessary number of training iterations for large-scale scenes, individual iterations take longer than standard 3DGS training due to data loading and hierarchy management overhead.
+Similarly, although rendering framerates are a significant improvement over neural-based and other out-of-core methods, they do not yet reach the peak performance of fully in-core 3DGS.
+Our approach requires roughly 1 GB of RAM per million Gaussians, which—while more efficient than prior methods—still constrains scalability for extremely large scenes.
+Loading from disk is feasible in our experiments, but at the cost of about a $10\times$ slowdown, making fast secondary storage highly desirable.
+The level-of-detail system makes our method robust to large variations in view distance.
+However, when such variation is absent—for example in single-height aerial datasets as evaluated in Suppl.
+A.3—the LoD machinery introduces unnecessary overhead, making more straightforward training a competitive alternative.
+Interactive rendering performance could be further improved by avoiding per-frame hierarchy cut recomputation and asynchronous streaming.
+While frustum culling effectively reduces memory load in most views, it is ineffective when the entire scene lies inside the frustum.
+Occlusion culling could address this limitation by skipping entire SPTs before they are loaded into memory.
+Overall, we believe that out-of-core 3D Gaussian Splatting is a promising direction for scaling radiance field methods to city-scale scenes and beyond on consumer-grade hardware.
+Appendix A Supplementary Material
+Our full supplementary material consists of:
+  1. (1)
+    This document, which includes additional ablation studies, implementation details and experiments, which we could not include in the main paper due to page restrictions
+  2. (2)
+    A short video presentation about this manuscript
+  3. (3)
+    An html-page with more qualitative comparisons and short videos
+  4. (4)
+    Our full source code contained in the Code directory
+  5. (5)
+    Training configurations detailing all hyperparameters for our and baseline methods on each experiment
+A.1. Additional Ablations
+We present additional experiments to demonstrate the effectiveness of each component of our method.
 View Selection
-When training MC-smaller-city+ dataset (250k iterations, up to 60M Gaussians), enabling our guided view selection reduces the number of Gaussians loaded from RAM per frame from 355 770 to 219 043—a $35\%$ reduction in memory transfers.
-Reconstruction quality is unaffected, with PSNR and SSIM improving slightly by 0.04 and 0.05, respectively.
+To verify the effectiveness of caching and view selection, we conduct an experiment by training the MC-smaller-city+ dataset (100k iterations, up to 30M Gaussians) with and without view selection and caching.
+The results in Table 4 show that while quality metrics are barely affected (with a maximal difference of 0.006 PSNR), the number of Gaussians that need to be loaded into VRAM decrease by a factor of almost $20\times$ with view selection and caching enabled.
+Table 4. Ablation results on the MC-smaller-city+ dataset. #Loaded refers to the average number of Gaussians loaded from RAM over 100 000 iterations.
+[…]
+w/o caching 21.831 0.701 0.425 1 629 187
+w/ caching 21.829 0.700 0.429 290 572
+w/ view selection 21.825 0.697 0.432 89 451
 Effectiveness of LoD
-To evaluate the effectiveness of our level-of-detail system, we choose 5 random images from each dataset and compare the resulting quality metrics and number of Gaussians rendered at 50 different levels of details.
-The results are visualized in Figure 18.
+To evaluate the effectiveness of our level-of-detail system, we choose 5 random images from each dataset and compare the resulting quality metrics to the number of Gaussians rendered at 50 different levels of details.
+The results are visualized in Figure 16 in addition to average FPS for rendering the view 10 times with pre-warmed cache on an H200 GPU.
+As expected, FPS scales closely with the inverse of the number of Gaussians.
 In the vast majority of cases, increasing the LoD level consistently improves rendering quality.
-Overall, we only notice a very small dip in quality when reducing the Gaussian count by 50%, verifying the effectiveness of our continuous LoD structure.
-(a) MC-small-city+ aerial
-(b) MC-small-city+ street
-(c) campus
-Figure 18: Quantitative evaluation of our LoD system for 5 randomly chosen views and 50 different levels of detail.
-A.4 Dataset Details and Further Results
+Overall, we only notice a small dip in quality when reducing the Gaussian count by 50%, verifying the effectiveness of our LoD structure.
+Further, using continuous instead of discrete levels of detail leads to the smoothness of the curves in Figure 16, indicating a smooth transition between LoD levels, which reduces popping.
+Effectiveness of Caching
+In Figure 15, we provide sensitivity curves for cache size vs.
+FPS during rendering.
+In each iteration up to cache size Gaussians from the previous frames may be reused.
+The results once again demonstrate the importance of the caching system to rendering efficiency.
+The graphs show that performance increases steeply with increasing cache size until it exceeds the typical number of Gaussians required per iteration and starts to plateau.
+(a) MC-small-city+
+(b) Campus
+Figure 15. Average FPS for rendering 1000 frames of the camera paths for the MC-small-city+ and Campus scenes with various cache sizes.
+(a) MC-small-city+ aerial.
+(b) MC-small-city+ street.
+(c) Campus.
+Figure 16. Qualitative evaluation of our LoD system for 5 randomly chosen views and 50 different levels of detail.
+A.2. Initialization and Training Details
+Following Kerbl et al.
+(2024), we initialize the Gaussian model from a sparse point cloud, augmented with skybox points.
+This initial representation is small enough to fully reside in GPU memory and is trained for 100100k iterations without densification.
+The goal of this phase is to establish a stable global scene structure before constructing the hierarchy.
+After this initial optimization, we build a binary Gaussian hierarchy, where the trained Gaussians act as leaf nodes and the parent nodes represent merged approximations of their two children.
+The Gaussian properties–including base color, SH coefficients, position, and covariance–are optimized using the standard loss propagation introduced by Kerbl et al.
+(2023).
+To fulfill our tight memory requirements, we update only the parameters of the cut set chosen for each training view.
+Thus, gradients may, in general, be propagated to Gaussians in the middle of the hierarchy.
+Assuming good coverage and diversity of training views, these updates will diffuse over the entire hierarchy over the coarse of training, leading to smooth transitions from the highest to the lowest LoD.
+A.3. Dataset Details and Further Results
 We include more details on the datasets used in the evaluation and provide additional results for smaller-scale scenes.
 MC-small-city+
-For our main benchmark, MC-small-city+, we aggregate 33 006 street-view images and 7 672 aerial views from the small-city scene of the MatrixCity dataset (MatrixCityDataset) and add 533 additional high-altitude views we generated manually.
+For our main benchmark, MC-small-city+, we aggregate 33 006 street-view images and 7 672 aerial views from the small-city scene of the MatrixCity dataset (Li et al., 2023) and generate 533 additional high-altitude views.
 We evaluate reconstruction quality on a separate set of 4 228 test views, according to the test split provided by the dataset.
-The scene is extremely challenging due to its enormous scale, sparse views and wide variation in scale: As such, we use a subset of the entire dataset (15.1k images, covering about a third of the area), denoted as MC-smaller-city+ for baselines that are not able to reconstruct the full dataset.
+The scene is extremely challenging due to its enormous scale, sparse views and wide variation in scale: As such, we also construct a subset of the entire dataset (15.1k images, covering about a third of the area), denoted as MC-smaller-city+ for baselines that are not able to reconstruct the full dataset.
 For MC-small-city+, we use the camera poses provided by the MatrixCity dataset and convert them to the COLMAP format for Gaussian splatting.
@@
 We do not make use of the ground truth depth images provided by the dataset, as we consider this to be an unrealistic advantage.
-Because the dataset does not provide point correspondences required for scaling monocular depth maps, we deactivate depth supervision in our method and baselines for this dataset.
 Figure 17 shows the scene along with the camera distribution.
-While some methods have successfully reconstructed only aerial (CityGaussian; CityGaussianV2; OctreeGS) or only street-level views (CityGaussianV2; NerfXL), training a model that holds up to scrutiny from both perspectives presents a particular challenge.
-Training on close and far views simultaneously–without a proper LoD system like ours in place–significantly degrades visual quality for both, as noted by HorizonGS; CrossView.
-Moreover, such a scenario complicates partitioning the scene into independent chunks.
-Figure 19: The MC-small-city+ dataset spans multiple city blocks, which are supervised by tens of thousands of views, varying widely in scale.
+While some methods have successfully reconstructed only aerial (Liu et al., 2024a, b; Ren et al., 2024) or only street-level views (Liu et al., 2024b; Li et al., 2024b), training a model that holds up to scrutiny from both perspectives presents a particular challenge.
+Training on close and far views simultaneously–without a proper LoD system like ours in place–significantly degrades visual quality for both, as noted by Jiang et al.
+(2025); Zhang et al.
+(2024).
+Moreover, such a scenario greatly complicates partitioning the scene into independent chunks.
+Why are the quality metrics on MC-small-city+ so low?
+When comparing our evaluation on the MatrixCity dataset with those in other works like Liu et al.
+(2024a); Zhao et al.
+(2026); Jiang et al.
+(2025); Ren et al.
+(2024), it might seem surprising that they achieve significantly lower quality metrics on our constructed MC-small-city+ scene.
+There are several reasons for this:
+  1. (1)
+    These works only evaluate on the aerial views, which make up less than 20% of the scene images and are generally easier to reconstruct.
+  2. (2)
+    Reconstructing street, aerial and the newly generated high-aerial views in the same scene is much more difficult than reconstructing just aerial views.
+  3. (3)
+    Some competing works (Liu et al., 2024a; Zhao et al., 2026) use the ground-truth depth maps or dense point clouds from the MatrixCity dataset.
+We would consider this to be an unrealistic advantage, since they contain almost the entire ground truth geometry of the scene.
+Zhao et al.
+(2026) even starts the reconstruction from the full dense point cloud and never densifies, whereas our evaluation starts from the same point cloud downsampled by a factor of 5.
+Figure 17. The MC-small-city+ dataset spans hundreds of buildings, which are supervised by tens of thousands of views, varying widely in scale.
+Uni10K
+In order to address the gap of real-world large-scale datasets that are captured across multiple scales, we present Uni10k.
+The scene consists of an outdoor campus of approximately $100\,000\text{ m}^{2}$, captured from both ground-level and aerial perspectives.
+Standard reconstruction via COLMAP at this scale would typically require weeks of computation due to the complexity of the image-matching and mapping stages.
+To mitigate this, we leverage spatial and temporal priors in conjunction with a coarse-to-fine scheme.
+Specifically, we utilize GPS data to reduce the matching complexity from quadratic to near-linear by limiting image comparisons to a predefined spatial radius.
+Regarding temporal priors, since frames are sampled from video sequences, we initially reconstruct a baseline model starting from a set of frames uniformly sampled every second of video.
+Then, we densify the camera coverage by incrementally registering, triangulating, and refining new images using only local BA.
+The process concludes with several rounds of Global bundle adjustment.
+The final reconstruction comprises more than 10 00010\,000 images at 4K resolution and  6.2M sparse points, with an overall MRE of approximately $0.63$ pixels.
+We train on the full resolution images and hold back every 8th frame alphabetically for the test set.
+Figure 18 shows the sparse point cloud and camera distribution.
+Figure 18. Uni10K’s sparse point cloud. The scene is reconstructed from more than $10$ thousand 4k images at widely varying heights, both from aerial and ground perspective.
 Hierarchical 3DGS dataset
-H-3DGS Hierarchical did not use a test split for their dataset, but we hold back every $100$th image alphabetically for testing.
+Since H-3DGS Kerbl et al.
+(2024) did not introduce a test split for their dataset, we hold back every $100$th image alphabetically for testing.
 We reevaluate H-3DGS for this scene with the new test split, using the camera calibrations and chunk splits provided on their website.
 In accordance with the instructions, we disable exposure optimization for evaluation.
-On this dataset, we use depth supervision identical to Hierarchical.
+On this dataset only, we use depth supervision identical to Kerbl et al.
+(2024).
+The distribution of street views in the Campus scene is visualized in Figure 19(b).
 Hierarchical 3DGS single-chunk dataset
-To demonstrate our ability to reconstruct small scenes, we also evaluate our method on the smaller, single-chunk versions of the Campus and Small City (cf.
+To demonstrate our ability to reconstruct small scenes, we also evaluate our method on the smaller, single-chunk versions of the Campus and Small City scenes (cf.
 Table 5).
 The quality metrics indicate that training on the LoD structure (which would not be required for scenes of this scales) during training only leads to marginal reductions in visual quality.
-The distribution of street views in the campus dataset is visualized in Figure 20(b).
-Table 4: H-3DGS single chunk view synthesis results. Results reproduced from Hierarchical.
+Table 5. H-3DGS single chunk view synthesis results. Results with are taken from (Kerbl et al., 2024).
 Small City Chunk (1.1k images) Campus Chunk (1.5k images)
 […]
+Mip-NeRF 360 24.70 0.765 0.348 20.95 0.731 0.442
+INGP-big 23.47 0.715 0.426 20.37 0.700 0.476
+F2-NeRF-big 24.53 0.762 0.342 19.25 0.681 0.478
 3DGS 25.34 0.776 0.337 23.87 0.785 0.378
-H-3DGS (leaves) 26.62 0.820 0.259 24.61 0.807 0.331
+H-3DGS 26.62 0.820 0.259 24.61 0.807 0.331
 Ours 25.94 0.808 0.276 24.86 0.796 0.360
-Mill19 dataset
-Table 5 shows results on the Mill19 dataset (MegaNerf).
-We use the camera poses and sparse point cloud provided by CityGaussian.
+OccluGaussian Dataset
+The OccluGaussian dataset features three large-scale indoor environments.
+Unfortunately, the download link for the Gallery scene on the official website leads to a broken zip file.
+As we have not received an answer from the authors, we were unable to evaluate on this scene.
+Additionally, some training images that are referenced in the camera pose files were not included in the download of the Canteen scene.
+We ran the evaluation on all images that were provided using full resolution images.
+Mill19 and Urbanscene3D dataset
+Table 6 and 7 show results on the Mill19 (Turki et al., 2022) and UrbanScene3D (Lin et al., 2022) datasets.
+We use the camera poses and sparse point cloud provided by Liu et al.
+(2024a).
 In accordance with baselines, we downscale all images by a factor of 4.
-We have chosen this dataset, because it is widely used in large-scale novel-view synthesis.
-At the same time, this dataset represents a worst-case scenario for our method, as the regular, same-height aerial views (cf.
+We have included these dataset, because they are widely used in large-scale novel-view synthesis.
+At the same time, they represents a worst-case scenario for our method, as the regular, same-height aerial views (cf.
 Figure 19(a)) negate any benefit of our level-of-detail method and can be split into independent chunks trivially.
-Table 5: Mill19 novel view synthesis results Results with are reproduced from CityGaussian. Results of H-3DGS are reproduced from Hierarchical.
+Table 6. Mill19 novel view synthesis results Results with are taken from Liu et al. (2024a). Results of H-3DGS are taken from Kerbl et al. (2024).
 Rubble (1.6k images) Building (1.9k images)
@@
 Ours 23.75 0.704 0.330 25M 20.67 0.682 0.318 25M
+Table 7. UrbanScene3D novel view synthesis Results. Results with are taken from Liu et al. (2024a).
+Residence (2.5k images) Sci-Art (3k images)
+[…]
+CityGaussian 21.90 0.805 0.217 21.34 0.833 0.232
+3DGS 21.44 0.791 0.236 21.05 0.830 0.242
+H-3DGS 17.52 0.416 0.391 18.42 0.652 0.340
+OctreeGS 18.76 0.523 0.539 18.97 0.620 0.535
+Ours 20.25 0.705 0.319 20.70 0.794 0.272
 (a) Rubble (Mill19)
 (b) Campus (H-3DGS)
-Figure 20: Overview of the aerial and street datasets.
-A.4.1 Rendering Comparison
-To demonstrate the effectiveness of our LoD method, we compare VRAM usage and image quality to rendering the highest level of detail leaf Gaussians with both 3DGS (3DGS) and gsplat (Gsplat) using their default rasterization settings.
-We conducted comparisons across three different viewpoints to assess performance at varying levels of detail.
-As illustrated in Figure 21, our approach achieves visual fidelity comparable to the baselines while significantly reducing VRAM usage, highlighting the effectiveness of our LOD scheme.
-(a) Overview
-(b) Aerial View
-(c) Street View
-Figure 21: Comparision to standard 3DGS (3DGS) and GSplat (Gsplat) in terms of rendering quality and VRAM usage.
-A.5 Evaluation Details
-Finding baselines for the MC-small-city+ scene proved challenging, as the only methods we are aware of that have successfully trained on all fused aerial and street-level views–RetinaGS and NerfXL–both require 64 GPUs running in parallel.
-We were not able to meet these hardware requirements for evaluation and hold that these methods and similar multi-GPU works like Grendel are orthogonal to our method.
-H-3DGS (Hierarchical) has demonstrated the ability to reconstruct similar sized datasets, but its evaluation was restricted to street-level views.
-We choose Hierarchical and other popular large-scale Gaussian Splatting methods (CityGaussian; HorizonGS; OctreeGS) as baselines.
-Note that ”large-scale” is often used ambiguously in the context of novel-view synthesis, alternatively meaning reconstructing very complex scenes with large number of images or reconstructing novel views using far-away (i.e.
-aerial) input views: With the exception of Hierarchical, the biggest datasets these methods use in their papers consist of just 5620 images.
-In order to train MC-small-city+ with 42.2 thousand images, it was necessary to modify CityGaussian, HorizonGS and OctreeGS to load images from disk instead of caching them in RAM or VRAM.
-This slows down training, but does not affect the final results.
-Most large-scale 3DGS frameworks have complicated multi-stage training processes and are very sensitive to hyperparameters.
-We consider it a benefit of our training pipeline that the user is not confronted with tuning the error-prone partitioning process, being more akin to the original Gaussian Splatting training.
-In general, the hyperparameters suggested for large-scale scenes were used for baselines.
-If provided, we use configurations explicitly targeting the MatrixCity dataset, if it improves metrics over default configuration.
-Learning rates were not modified, as they were similar to ours and should not be a barrier to reconstruction on this scale.
-We also reached out to the authors of all our baselines about suggesting hyperparameters for the MC-small-city+ dataset, which we incorporated when provided.
-For methods that make use of configuration files, we include them in the supplemental material.
+Figure 19. Overview of the aerial and street datasets.
+A.4. Evaluation Details
+Finding baselines for the MC-small-city+ scene proved challenging, as the only methods we are aware of that have successfully trained on all fused aerial and street-level views–Li et al.
+(2024a) and Li et al.
+(2024b)–both require 64 GPUs running in parallel.
+We were not able to meet these hardware requirements for evaluation and hold that these methods and similar multi-GPU works like Zhao et al.
+(2024) are orthogonal to our method.
+H-3DGS (Kerbl et al., 2024) has demonstrated the ability to reconstruct similar sized datasets, but its evaluation was restricted to street-level views.
+We choose Kerbl et al.
+(2024) and other popular large-scale Gaussian Splatting methods (Liu et al., 2024a; Jiang et al., 2025; Ren et al., 2024; Zhao et al., 2026) as baselines.
+Table 8 compares the supported features of our method and all baselines.
+Most of the large-scale 3DGS frameworks have complicated multi-stage training processes and are very sensitive to hyperparameters, which we detail in the following section.
+We consider it a benefit of our training pipeline that the user is not confronted with tuning the error-prone partitioning process, being more similar to the original Gaussian Splatting training.
+LoD Training LoD Rendering Out-of-Core Training Interactive Rendering No Chunking Artifacts Gaussians
+CityGaussian Explicit
+HorizonGS Neural
+H-3DGS Explicit
+OctreeGS Neural
+CLM-GS Explicit
+VastGaussian Explicit
+OccluGaussian Explicit
+Ours Explicit
+Table 8. Comparison of features.
 OctreeGS
-For OctreeGS (OctreeGS), we follow their instructions on training custom datasets.
-Note that OctreeGS does not perform any scene division and uses the memory-intensive ScaffoldGS (ScaffoldGS) method, making it unsuitable for scenes of this scale.
+For OctreeGS (Ren et al., 2024), we follow the provided instructions on training custom datasets.
+We trained using the suggested hyperparameters for standard scenes and those used for the MatrixCity dataset.
+The latter achieved higher quality metrics, which is what we report.
+Note that OctreeGS does not perform any scene division and uses the memory-intensive ScaffoldGS (Lu et al., 2024) method, making it unsuitable for ultra-large scale scenes.
 CityGaussian
-We choose CityGaussian (CityGaussian) (specifically the 1.2 version of the repository) as a baseline instead of CityGaussianV2 (CityGaussianV2), as the code release for CityGaussianV2 does not yet support level-of-detail rendering.
+We choose CityGaussian (Liu et al., 2024a) (specifically the 1.2 version of the repository) as a baseline instead of CityGaussianV2 (Liu et al., 2024b), as the code release for CityGaussianV2 does not yet support level-of-detail rendering.
 We follow the instructions on training large datasets and use the parameters and chunk split from the provided configuration file for the MatrixCity-Aerial dataset (which covers the same region as the MC-small-city+ dataset, but without the street views), as they produced better metrics than the default parameters.
 Note that the low number of Gaussians in the final model is due to CityGS discarding most (about 90%) of the trained Gaussians after chunk training in order to avoid artifacts.
-For campus and small city, we used a $4\times 4$ and $2\times 2$ chunk split to match the number of chunks as closely as possible with the 1212/44 chunks used by H-3DGS (Hierarchical).
+For Campus and Small City, we used the same parameters, but a $4\times 4$ and $2\times 2$ chunk split to match the number of chunks as closely as possible with the $12$ and $4$ chunks respectively used by H-3DGS (Kerbl et al., 2024) and a $3\times 3$ chunk split for the Uni10k dataset.
+CLM-GS
+Significant modifications were necessary to evaluate CLM-GS (Zhao et al., 2026) on our large-scale datasets.
+In the paper, the issue of large-scale densification is sidestepped, by starting from a dense (102 million point) ground truth point cloud and disabling densification all-together for the aerial matrix-city scene.
+All other evaluated scenes are sufficiently small that densification is not an issue.
+Running the unmodified method on large-scale datasets from sparse point clouds leads to either collapsing training or pruning of all Gaussians.
+Thus, we disable pruning to avoid large parts of the scene disappearing and disable opacity resets, which make the training process unstable for large-scale scenes.
+Experimentally we found that the high densification parameters for the 28M mode of Rubble produced the best results and used it in our evaluation.
+We further implemented the suggestion of Kerbl et al.
+(2024) to densify according to maximal instead of average screen-space gradients, but did not notice significant improvements.
+As suggested by the authors for large datasets, we use the clm-offload strategy with a batch size of 4, and match the number of training iterations with our method for each experiment.
+We enable the option to load images from disk in order to avoid going out of memory.
+As of writing the CLM codebase contains no possibility of rendering test images or evaluating performance metrics, so we implemented the necessary functions ourselves.
 Hierarchical-3DGS
-We follow the instructions of Hierarchical for running the method on large scenes.
+We follow the instructions of Kerbl et al.
+(2024) for running the method on large scenes.
 Note that we do not make use of their particular COLMAP pipeline for MC-smaller-city+, as it only works with unbroken camera paths.
@@
 On scenes of this scale, COLMAP output contains significant noise, which also leads to reduced quality metrics for our method.
-H-3DGS achieves drastically worse PSNR and SSIM results on the full campus dataset, compared to the single-chunk results, due to a slight perspective distortion that occurs specifically on this dataset (it did not occur on small city).
-This can be reproduced by evaluating the trained model on their campus dataset, both of which are available on their website.
-As such, we include the results, but encourage visual comparison in Figure 6.
+H-3DGS achieves drastically worse PSNR and SSIM results on the full Campus scene, compared to the single-chunk results, due to a slight perspective distortion that occurs specifically on this dataset (it did not occur on Small City).
+This can be reproduced by evaluating the trained model on their Campus scene, both of which are available on their website.
+As such, we include the results, but encourage visual comparison in Figure 12 of the main paper.
 HorizonGS
-We expand the configuration provided for a single chunk of the MatrixCity small-scene dataset (named ours/large_scene/block_A) with a chunk split of $3\times 4$ and $5\times 5$ for MC-smaller-city+ and MC-small-city+ (shown in Figure 22) respectively.
-The partitions of the MC-small-city+ scene Each chunk is trained for 60000 iterations and then fine-tuned for another 40000.
-Figure 22: Partitions of HorizonGS on the MC-small-city+ scene.
+We expand the configuration provided for a single chunk of the MatrixCity small-scene dataset (named
+ours/large_scene/block_A) with a chunk split of $3\times 4$ and $5\times 5$ for MC-smaller-city+ and MC-small-city+, respectively, and $2\times 2$ for Uni10k.
+Since the method throws an error if it is not given both a set of street- and aerial images, we limit evaluation to those two scenes.
+Each chunk is trained for 60000 iterations and then fine-tuned for another 40000.
+The chunk partitions are shown in Figure 20; as we can see, the complexity of these partitions is vastly different.
+For Uni10k, the simpler chunk partitioning leads to more stable training, and, as can be seen in Table 1, improved results.
+Figure 20. Partitions of HorizonGS on MC-small-city+ (left) and Uni10k (right).
 VastGaussian
-VastGaussian (VastGaussian) would have presented a natural comparison point, but there is currently no official code release available.
+VastGaussian (Lin et al., 2024) would have presented a natural comparison point, but there is currently no official code release available.
+OccluGaussian
+As of writing, the code for (Liu et al., 2025) has not been released, so we are only able to compare to their reported results on their own dataset.
+Because the method is specifically tailored to indoor scenes, they never published results on any outdoor street-level datasets.
 Ours
-We provide detailed configuration files for all of the scenes as supplemental material.
-In general, we perform 60k iterations of coarse training (except for 100k on MC-small-city+) and then 150k iterations of fine training (except 250k on campus and MC-smaller-city+ and 500k on MC-small-city+).
-We shoose SH degree 1 and half the densifaction gradient threshold as opposed to H-3DGS.
-We pick a cache size of 15000000 Gaussians.
-A.6 Implementation Details
+In general, we perform 60k iterations of coarse training (except for 100k on MC-small-city+) and then 150k iterations of fine training (except 100k on Uni10k, 250k on Campus and MC-smaller-city+ and 500k on MC-small-city+).
+We choose SH degree 1 and halve the densification gradient threshold as opposed to H-3DGS.
+We pick a cache size of 15 million Gaussians.
+For each scene, we pick a maximal number of Gaussians and desired SPT volume according to the scene scale.
+We deactivate the use of pinned CPU memory, which would accelerate memory transfers, because it would not be available in sufficient quantities on the consumer devices we are targeting.
+We provide detailed configuration files containing all hyperparameters for all of the scenes as supplemental material.
+A.5. Implementation Details
 Training models on this unprecedented scale on consumer hardware required solving many technical issues. In this section, we elaborate on the most important of these design decisions.
 Codebase
-Our general code structure is based on H-3DGS (Hierarchical).
-In particular, we reuse the hierarchy creator, but replace the cut procedure and chunk based training, as well as the rasterizer for the one used in 3DGS.
-We also use the densification implementation of 3DGSasMCMC.
+Our general code structure is based on H-3DGS (Kerbl et al., 2024).
+In particular, we reuse the hierarchy creator, but replace the cut procedure and chunk based training, as well as the rasterizer for the one used in Kerbl et al.
+(2023).
+We also use the densification implementation of Kheradmand et al.
+(2024).
 All of the code is implemented in PyTorch and C++/CUDA.
@@
 Therefore, we choose a base focal length $f_{b}$ and use a relative distance metric $\hat{d}$ for our HSPT cuts, which is calculated for the current camera’s focal length $f_{i}$ as $\hat{d}=\frac{f_{b}}{f_{i}}d$.
-E.g.
-for a camera with double the focal length, the cut distance should be halved to account for the larger projection footprint on the image plane.
+For example, given a camera with double the focal length, the cut distance should be halved to account for the larger projection footprint on the image plane.
 SH Degree
@@
 The properties of all required Gaussians are then transferred to the GPU via a single copy operation, reorganized on the GPU in a structure of arrays layout (as required by the rasterizer), and appended to the current render set.
-A.7 Additional Densification Details
-Our densification method combines the split operation from 3DGSasMCMC with the selection method from Hierarchical.
-Relying purely on the densification strategy of 3DGSasMCMC–applying a noise to every Gaussian’s position and an opacity/scaling loss–becomes problematic for large scenes, where a majority of Gaussians are not visible from any one view: The result is a gradual disappearance of density in the scene or extreme ”stringing” artifacts (cf.
+A.6. Additional Densification Details
+Our densification method combines the split operation from Kheradmand et al.
+(2024) and the selection method from Kerbl et al.
+(2024) with coarse-to-fine expansion of the LoD structure.
+Relying purely on the densification strategy of Kheradmand et al.
+(2024)–applying a noise to every Gaussian’s position and an opacity/scaling loss–becomes problematic for large scenes, where a majority of Gaussians are not visible from any one view: The result is a gradual disappearance of density in the scene or extreme ”stringing” artifacts (cf.
 Figure 21).
@@
 On small scenes such as the single-chunk datasets, our modified MCMC strategy slightly outperforms the presented strategy given sufficient Gaussian budget, but it lags behind in the larger scenes.
+Avoiding destructive opacity resets, as advocated by Kerbl et al.
+(2023), is particularly important in our setting, as many Gaussians will no longer conform to the hierarchical structure after opacity is restored to normal.
 (a) Campus
 (b) Small City
-Figure 23: Without regularization, the scale- and opacity-losses of 3DGS-as-MCMC lead to either stringing or disappearance of the entire scene.
-A.8 Pseudocode
+Figure 21. Without regularization, the scale- and opacity-losses of 3DGS-as-MCMC lead to either stringing or disappearance of the entire scene.
+A.7. Pseudocode
 This section provides pseudocode for our core algorithms.
@@
 end procedure
-A.9 Supplemental Material
-As supplemental material, we include the codebase used for this project, along with the configuration files for all of our scenes.
-We also include configuration files for our baselines, if applicable.
-Further, we provide videos at the following anonymous link: https://www.youtube.com/playlist?list=PLocO6QqSKZ4YN7DyuC5gWEeVciSV6RdAk.
-This includes fly-throughs of the evaluated scenes and a method overview video.
-A.10 Declaration of LLM usage
-LLMs were used to polish the language in the paper, but did not write any sections on their own. Some code sections in the implementation were generated using LLMs.